### v4

# TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented Large Language Models

Abstract.
Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming. With these tools, developers can easily generate the desired complete functional code based on incomplete code snippets and natural language prompts. Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks. The former poisons the training data or model parameters, hijacking the LLMs to generate malicious code snippets when encountering the trigger. The latter crafts malicious adversarial input codes to reduce the quality of the generated codes. However, both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes.
This paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA). TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base). It induces the Code LLMs to generate attacker-specified malicious code snippets at the target location. In general, we compress the attacker-specified malicious objective into the perturbation by adversarial optimization based on greedy token search. We collect 13 representative malicious objectives to design 31 threat cases for three popular programming languages. We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerts a new practical threat of using Code LLMs.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation. In particular, code-oriented large language models (Code LLMs), e.g., CodeX (Wojciech Zaremba, 2021), CodeGemma (Team et al., 2024), and CodeGeex (Zheng et al., 2023), have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus (Zhang et al., 2023; Li et al., 2023b; Ding et al., 2024) and the fundamental paradigm in commercial practice (Friedman, 2021; cod, 2024).. They can efficiently generate complete functional code based on incomplete code snippets and text prompts. Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot (Friedman, 2021) developed jointly by Github (Git, 2024) and OpenAI (Ope, 2024). These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
Despite their remarkable performance, recent studies revealed that Code LLMs are vulnerable to some well-designed attacks, which maliciously manipulate the functional execution of Code LLMs in practice. In general, existing attacks can be divided into two main categories. (1) Backdoor attacks (Schuster et al., 2021; Li et al., 2023a; Yang et al., 2024). They intend to compromise the models by poisoning the training data (Schuster et al., 2021; Yang et al., 2024) or model parameters (Li et al., 2023a), such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns. They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications (Li et al., 2023a), or implementation of less secure encryption rules or protocols (Schuster et al., 2021). (2) Adversarial attacks (Zhou et al., 2022; Bielik and Vechev, 2020; Jha and Reddy, 2023). They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.(Jha and Reddy, 2023) slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score (Ren et al., 2020).
However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent. Specifically, backdoor attacks can achieve specific malicious objectives but require the capability of interfering the model training, e.g., modifying some training samples or even model parameters. In real-world scenarios, developers of Code LLMs or integrated applications usually obtain the models or datasets from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks (Bielik and Vechev, 2020; Zhou et al., 2022; Jha and Reddy, 2023) can hardly achieve precise malicious objectives, although they do not require intervention in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input. Driven by these findings, an intriguing and critical question arises: Has the application of Code LLMs already achieved a sufficient level of security in practice?
Unfortunately, the answer to the above question is negative. In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger (Wallace et al., 2019) and Greedy Coordinate Gradient (GCG) (Zou et al., 2023). Please find more details in Section 2 and Appendix A.1.. TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks. In general, as shown in Figure 1, the TPIA attacker generates a concise non-functional perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion) , and inserts it into the victim’s code context. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations. Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice. For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs (Chen et al., 2024) via retrieval-augmented generation (Lewis et al., 2020). In particular, different from supply chain attacks (Ohm et al., 2020) that can be easily detected by static analysis (Li et al., 2018b) or sandboxing (Alhamdan and Staicu, 2023) since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
Figure 1. The attack paragdim of our TPIA. A concise non-functional attacker-generated perturbation is injected into one of the project’s files. When the victim user completes the code via the Code LLM, the model will be induced by the incorporated attacker-specified code to generate an unsafe output at specified position, e.g., in edit_part() function.
In general, our attack consists of two main stages. In the first stage, we construct the input and target output based on the attacker-specified objective. We represent the malicious objective as an output tuple (target position, target code), which aims to induce the Code LLM to generate the target code at the target position. We further expand the tuple into a complete task code, including a conditional code (e.g., necessary package references), context code (e.g., noises), and position code to simulate the actual attack scenario. In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective. Specifically, we obtain the base Code LLM of the integrated application and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code. This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. (2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack.
We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) (Top, 2024) across 3 programming languages. There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++. We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b, and CodeGeex2-6b. The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness. In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX (cod, 2024) (with more than 535k installations in the VScode (VSC, 2024) extension shop) and Github Copilot (Friedman, 2021) (developed by Github and OpenAI (Ope, 2024)). These applications correspond to the gray-box and black-box attack scenarios, respectively. Our TPIA achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its practical threat in the real world.
In summary, our main contributions are three-fold.
  • 1)
    We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs.
  • 2)
    We introduce a new attack paradigm (i.e., TPIA) against Code LLMs. It can induce the Code LLM to generate an attacker-specified malicious code snippet at specific locations by simply inserting a concise perturbation into the victim’s code context.
  • 3)
    We conduct extensive experiments to evaluate our TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 widely used commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
2. Background and Related Work
2.1. Code LLMs for Code Completion
Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation (Sun et al., 2022; Roziere et al., 2020), and code summarization (McBurney and McMillan, 2016; Ahmad et al., 2020). Many models have been developed to achieve these tasks, including CodeGen (Nijkamp et al., 2022), CodeBERT (Feng et al., 2020), and CodeT5 (Wang et al., 2021). The emergence of LLMs provides new alternatives to the coding tasks. By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex (Wojciech Zaremba, 2021), CodeLlama (Roziere et al., 2023), CodeGemma (Team et al., 2024) and CodeGeex (Zheng et al., 2023). These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast array of functionally complex code (Xia et al., 2023).
In this paper, we focus on code completion (Li et al., 2023b; Ding et al., 2024; Zhang et al., 2023). This is a paradigmatic task that involves generating new code from incomplete input (with natural language) using Code LLMs, which have been integrated into commercial applications (Friedman, 2021; cod, 2024). However, this development paradigm may encounter some potential threats. On the one hand, Code LLMs can be easily influenced by the code context, which alters the result of code generation. On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs. These threats have already occurred in the real world (Cri, 2025), highlighting the necessity of its security research.
2.2. Backdoor and Adversarial Attacks
Backdoor attacks target the training process of DNNs (Liu et al., 2020; Li et al., 2022; Zhang et al., 2024). These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present. In the code domain, there are generally two strategies to achieve this, including dataset poisoning (Schuster et al., 2021; Yang et al., 2024) and model poisoning (Li et al., 2023a). The former requires the attacker to modify the code dataset used for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM used for application development. However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned research institutes or companies). This mitigates the backdoor threats from the source. Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult. Arguably, backdoors cannot be implanted into a Code LLM without controlling the model training process.
Different from backdoor attacks, adversarial attacks primarily target the model’s inference process (Li et al., 2018a; Doan et al., 2021; Wang et al., 2022). Adversarial attacks induce the vulnerabilities of models by introducing well-crafted perturbations to the input data, leading the model to make incorrect output. In the code domain, existing adversarial attacks (Bielik and Vechev, 2020; Zhou et al., 2022; Jha and Reddy, 2023) aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs. For example, Jha et al.(Jha and Reddy, 2023) slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score (Ren et al., 2020). However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process. This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input. Accordingly, adversarial attacks are usually difficult to pose a serious threat to Code LLMs in practice.
2.3. Prompt Injection and Indirect Prompt Injection Attacks
Prompt injection (PI) attacks, are a special form of attack targeting LLMs (DAN, 2023; Liu et al., 2023b; Zou et al., 2023). A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb. A general way to carry out these attacks is by submitting a carefully designed prompt query to the LLM. In the code domain, existing research (Jiang et al., 2023; Liu et al., 2023a; Fang et al., 2024) focuses on code-oriented attack objectives, such as using LLMs to create malware. In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
Indirect prompt injection (IPI) attack (Greshake et al., 2023), on the other hand, considers a different attack scenario than PI. The attacker attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites or providing false information. This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses). In the code domain, the only existing work (Greshake et al., 2023) makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it needs to insert a large amount of readable comments containing the malicious intent, which can be easily noticed by programmers, as shown in Figure 11. Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
3. Problem Statement
3.1. Code Completion Task
Code completion is an essential task for Code LLMs. The input to the models may include code snippets, comments, pseudocode, and natural language requests. The output should consist of executable code blocks that meet the requirements specified by the input. In this paper, we consider a general task format. Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},x_{3},...,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements. For instance, natural language requests can only be included in the code comments. The output is the completed code block $\bm{Y}=\{x_{1},x_{2},x_{3},...,x_{m},x_{m+1},x_{m+2},...,x_{n}\}$, which should be logically executable and functionally compliant with the requirements. We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
3.2. Target-specific Attack
In the code domain, target-specific attack refers to the scenario where the attacker makes the Code LLM generate unsafe code for specific targets. Generally, the attacker’s targets can be described as a tuple of two continuous token sequences $(\bm{Y_{T}},\bm{T})$, where $\bm{Y_{T}}=\{x^{t}_{1},x^{t}_{2},...,x^{t}_{j}\}$ is a subsequence within $\bm{Y}$, and $\bm{T}=\{t_{1},t_{2},t_{3},...,t_{k}\}$ is what the attacker aims to insert after $\bm{Y_{T}}$ to create malicious functionality. To achieve the targets, the attacker needs to design a perturbation $\bm{P}$, represented as $\bm{P}=F(\bm{Y_{T}},\bm{T})$, and insert it into the input. The function $F$ can be implemented as manually creating the perturbation or by an optimization-based method as proposed in this paper. It is worth noting that, unlike backdoor attacks, the connection between $\bm{Y_{T}}$ and $\bm{T}$ is not built during the training process. To conclude, for a successful target-specific attack, the model output received by the victim $\bm{Y_{v}}$ will be:
(1) $\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\}.\end{split}$
3.3. Threat Model
Attack Scenarios. We consider common scenarios where users introduce external code by copying or installing packages. The attacker generates a concise non-functional perturbation containing malicious instructions and inserts it into the victim’s code context via the external code. When the user requests the Code LLM to perform code completion based on code context, the perturbation will induce the model to generate an unsafe output at the specified location. In particular, different from supply chain attacks (Ohm et al., 2020) that can be easily detected by existing main-stream detection methods (e.g., static analysis (Li et al., 2018b)) since they directly insert malicious functional code, to implant malicious behaviors, our attack scenario is more threatening because the attacker can circumvent detection methods by introducing only non-functional malicious code. Besides introducing code to the victim’s local project, attackers can upload their malicious payload codes to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs (Chen et al., 2024) via retrieval-augmented generation (Lewis et al., 2020). Code in the knowledge base can also appear in the context of the Code LLM’s prompts.
Attacker’s Capabilities. We assume that the attacker can obtain the parameters of Code LLM to generate malicious perturbations. Arguably, when developers build integrated applications using open-source models (such as CodeGemma (Team et al., 2024), CodeLlama (Roziere et al., 2023), or CodeGeeX (Zheng et al., 2023)), obtaining the corresponding model parameters is feasible. In this scenario, TPIA can achieve optimal performance. In this paper, we also consider gray-box (i.e., with unknown model version and system prompts) and black-box (i.e., with closed-source commercial models) scenarios, based on two widely-used commercial Code LLM-integrated applications in Section 5.3. We conduct successful attacks against these applications via perturbations generated with a local open-source Code LLM, confirming the practical threat of our TPIA in the real world.
On the other hand, the attacker needs to submit the payload code into the code repository and prevent it from being detected. Since our threat model uses non-functional code to achieve objectives, the payload code does not need to be designed with malicious functionality. In other words, the attacker can freely design the functional part of the payload code to ensure successful submission without worrying about detection by methods based on malicious functionality. In contrast, the perturbation part of the payload needs to be stealthy enough to avoid detection by victim users.
Attacker’s Goals. The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code). This tuple follows the definition of target-specific attacks in Section 3.2. Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$. The attacker attempts to insert a piece of designed malicious code $\bm{T}$ after a segment of the victim’s normal code $\bm{Y_{T}}$, and cause the victim to receive abnormal output code $\bm{Y_{v}}$. To achieve this goal, the attacker applies the adversarial methods to compute a perturbation $\bm{P}$ from the Code LLM’s parameters $\bm{\theta}$, which is most likely to induce the Code LLM to complete $\bm{T}$ after $\bm{Y_{T}}$. It can be formulated as below:
(2) $\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$
Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section 3.2 and the generated perturbation $\bm{P}$ should satisfy Equation (1).
In particular, the attacker can achieve these goals by inserting code snippets into specified positions, without needing to oppose the Code LLMs’ safety alignment. For example, when setting the target position to DES.new( and the target code snippet to key,DES.MODE_ECB), the generated perturbation can induce Code LLMs to use the easily cracked ECB mode for encryption; When setting the target position to TRG() (a shorthand for Target which can be replaced with any function name), the generated perturbation can induce Code LLMs to insert designed code snippet into the TRG function or after using it.
Figure 2. The overall pipeline of our TPIA attack. It is executed with two attack steps. (1) Objective Transfer. The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective. (2) Perturbation Injection. The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub (Git, 2024) and StackOverflow (Sta, 2024)) and introduced to the victim’s code context via various paths, such as direct copy and installing packages. When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
4. Methodology
4.1. Overview
We introduce our target-specific prompt injection attack (TPIA). In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2. Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section 4.2). In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section 4.3). The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section 4.4); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section 4.5). We describe their technical details in the following parts.
4.2. Task Construction
In the first stage, TPIA constructs specific code completion tasks based on the malicious objectives. The constructed tasks simulate the victim’s local code environment and are used to optimize the malicious perturbation, which is closely related to the final performance of TPIA. Each task can be divided into two parts: the input containing the target position, and the expected output containing the target code. This task is to be constructed so that our attack satisfies the following three properties to ensure effectiveness.
Property 1: Feasible Malicious Intent. The code environment configured by the user must meet the basic requirements for the malicious objective, including the necessary packages and methods declarations. Therefore, we introduce the conditional code at the beginning of the input code.
Property 2: Accurate Code LLMs’ Response. The response of the Code LLMs should be the target position. As stated in Section 3, the target code should be inserted after the target position to achieve the malicious objective. Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
Property 3: Robust Malicious Perturbation. The attacker cannot anticipate user’s contextual code in advance, which is the inevitable noise for TPIA. Therefore, we design the context code and add irrelevant noise to enhance the robustness of the malicious perturbation. The context code is placed between the position code and conditional code.
To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode, i.e., making the Code LLM insert Target Code (‘‘DES.MODE_ECB’’) after “cipher = DES.new(key,”. To achieve this, TPIA constructs an output tuple as (“cipher = DES.new(key,” , “DES.MODE_ECB)”. The other three task components are automatically generated with the following stages. Firstly, we declare the Crypto series methods that enable the encryption code to run normally in the conditional code. Secondly, we add “cipher = DES.new(key,” to the last line of the position code to ensure the continuity with the target output “DES.MODE_ECB)”. In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”. Finally, TPIA constructs the code unrelated to the task, such as “generate_random_string”, and uses it as the noise-like context code. Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
Figure 3. The example of task construction that induces Code LLM to use encryption mode with low security. In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
4.3. Malicious Perturbation Generation
In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code. Utilizing the code completion task constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer. Through a greedy search approach, the optimal perturbation is derived. The optimization process is illustrated in Algorithm 1, with detailed descriptions below.
Algorithm 1 Perturbation Tokens Optimization
1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{T}$, initial perturbation $\bm{P}$ with adversarial parts $x_{\text{adv},1:l}$, iterations $k$
2:  Vocab set $V$, $\mathbf{e}\in V$
3:  repeat
4:     for $i\in[0,\ldots,l]$ do
5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P},\bm{\theta})$
6:        $\mathcal{G}=\nabla_{x_{\text{adv},i}}\mathcal{L}$
7:        $\mathbf{x_{\text{candidates}}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
8:        for $j\in[0,\ldots,k]$ do
9:           $\bm{P}^{\prime}:x_{\text{adv},i}\to x_{\text{candidates}}^{j}$
10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P^{\prime}},\bm{\theta})$
11:           if $\mathcal{L}^{\prime}<\mathcal{L}$ then
12:              $\bm{P}=\bm{P}^{\prime}$
13:           end if
14:        end for
15:     end for
16:  until perturbation did not change in the last iteration
17:  Output: Optimized perturbation $\bm{P}$
Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation. Considering that our enhancement method adds static tokens to the perturbation (Section 4.5), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$. After determining the input and output of the code completion task, TPIA performs continuous iterations to optimize the perturbation. Specifically, since the attack is ultimately executed by the perturbation embedded in the external source code, in each iteration, TPIA only replaces the adversarial part of the perturbation token by token (Line 4). We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
(3) $\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$
where $\bm{T_{s}}$ denotes the task code. We only consider the prediction probability of each token in the target code when designing the loss function since the success of TPIA is solely related to whether the target code is correctly generated.
Besides, the optimization objective for the $i$-th adversarial token $x_{\text{adv}_{i}}$ can be denoted as:
(4) $\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$
To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer. Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
(5) $e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$
where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss. TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7). This has been proven to be highly effective in existing NLP tasks (Zou et al., 2023; Wallace et al., 2019). Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9). It then computes the new loss value using $\bm{P^{\prime}}$ and compares it with the original one: if the new value is better, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13). Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger (Wallace et al., 2019) and GCG (Zou et al., 2023). We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection). However, its effect is still very limited due to the difficulties of targeted malicious tasks. For example, as shown in Figure 9, its attack success rate (0-key, no FRE) is 0%. To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA. More details about the comparison between existing methods and TPIA are discussed in Appendix A.1.
Figure 4. Loss function distortion during the forward reasoning process. When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence. However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective.
4.4. Forward Reasoning Enhancement
This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness. Specifically, the calculation of the loss function in Malicious Perturbation Generation will encounter the distortion issue during the actual reasoning process of the Code LLMs. As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations. For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context. In our experiment, a malicious perturbation might have a much smaller loss value compared to a successful manually-crafted perturbation, yet the former fails to produce the desired output. This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions. This indicates that a deviation occurring during the sampling of preceding tokens can distort the probability calculations for subsequent tokens.
To alleviate this problem, we propose a forward reasoning enhancement technique. Our intuition is that the former tokens of the target code should be assigned more weights to enhance the forward reasoning. Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error. Its formulation is shown as follows:
(6) $\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$
$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
where $\mathcal{L}_{e}$ represents the additional reasoning-based enhancement applied to the original loss function, and $h$ is a hyperparameter set to either 1 or 2, which significantly improve the performance of attack in practice.
Figure 5. Examples of the keyword-based perturbation and other perturbations. The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage. Our keyword-based perturbation can balance both effectiveness and stealthiness.
4.5. Keyword-based Perturbation Design
This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack. Specifically, since the content of the perturbation and the payload code is not logically consistent, the attacker must enhance perturbation’s stealthiness as much as possible to avoid human detection. From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack. Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent. However, although we can use some optimization methods to obtain the optimal perturbation, it is still difficult to use illegible tokens to express complex instructions. A typical failure case is that even after the optimization, the attack objective cannot be completely transferred into the perturbation, resulting in attack invalidation. A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues. First, the malicious perturbation cannot gain any advantage in terms of the token count. Second, this leads to ineffective optimization efforts. As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds. Additionally, the complex tokens generated after adversarial computation often occupy more vision space.
To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design. This method allows the malicious perturbation to consume fewer tokens to convey malicious instruction, enhancing the attack stealthiness. Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens. Specifically, we need to use non-sensitive information that does not directly contain malicious features, which we call keywords, as static tokens for perturbation. For instance, as shown in Figure 5, part of the target function name (TR) is the key information, which is not directly related to the malicious objective (data deletion). The tokens of keywords are selected from the position code and target code after each perturbation optimization process, and the content is ensured to be non-malicious by limiting the number and position of selected tokens. The details of this method is discussed in Appendix A.2. By initializing the perturbation as “keywords<AT><AT><AT><AT>”, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
5. Evaluation
In this section, we conduct an extensive evaluation of TPIA, demonstrating its threat to Code LLMs from the following perspectives:
  1. 1)
    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section 5.2)
  2. 2)
    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section 5.3)
  3. 3)
    TPIA’s modules are all effective in improving attack performance. (Section 5.4)
5.1. Experiment Settings
Table 1. The summary of objectives and datasets. We list the programming languages that each objective/dataset is applicable: C/C++ (C), Python (P), Java (J). The last row of the table shows the total number of test data we design. Some objectives apply only to C/C++, resulting in a total of 31 applicable cases.
Identifier/Name Applicable Languages
Objectives ST0, ST1, ST2 C/C++, Python, Java
CWE416, 122, 787, 843 C/C++
CWE20, 78, 502, 918, 22, 306 C/C++, Java
Datasets MBPP Python
Humaneval Python
Eval-Plus Python
CodeXGLUE Python,Java
Humaneval-x C/C++,Python,Java
Data Count: 14273 (C/P/J): 2132/7965/4176
Table 2. The ASR across different Code LLMs. The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs. As shown in Table 1, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets. Best attack results are marked in bold.
[…]
  CodeGemma-2b No Attack 1.9% 17.3% 9.4% 22.0% 4.9% 12.6% 17.0% 12.2%
C-GCG (Zou et al., 2023) 40.6% 50.2% 32.1% 36.5% 49.4% 59.0% 44.2% 44.6%
IPI-M (Greshake et al., 2023) 3.4% 30.8% 6.4% 33.3% 3.2% 19.7% 31.0% 18.3%
TPIA (ours) 75.5% 93.4% 84.2% 74.4% 67.3% 94.1% 85.4% 82.0%
CodeGemma-7b No Attack 0.0% 12.2% 0.0% 18.9% 0.0% 15.6% 12.1% 8.4%
C-GCG (Zou et al., 2023) 43.5% 53.8% 55.4% 33.1% 20.4% 47.2% 50.6% 43.4%
IPI-M (Greshake et al., 2023) 11.1% 29.6% 4.3% 46.8% 20.8% 37.8% 33.2% 26.2%
TPIA (ours) 85.8% 97.9% 78.5% 81.6% 70.5% 94.4% 97.4% 86.6%
CodeGeeX2-6b No Attack 0.1% 13.4% 12.9% 19.0% 24.0% 21.9% 14.3% 15.0%
C-GCG (Zou et al., 2023) 42.9% 46.1% 44.9% 41.7% 32.5% 51.7% 33.7% 41.9%
IPI-M (Greshake et al., 2023) 9.5% 11.4% 5.3% 17.9% 0.9% 22.0% 11.5% 11.2%
TPIA (ours) 79.6% 95.0% 80.8% 83.3% 72.0% 97.9% 85.6% 84.9%
Malicious Objectives and Code Datasets. We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) (Top, 2024) across 3 programming languages. The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’). ST0 and ST1 are studied in classical code backdoor attacks (Schuster et al., 2021), while ST2 is a famous destructive instruction. CWE vulnerabilities are generally selected for experiments in code model research (He and Vechev, 2023; Yan et al., 2024), and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA). We consider three mainstream programming languages: Python, Java, and C/C++ to design the threat case for each objective. All the cases are constructed by the method described in Section 4.2 based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2). We will discuss the cases in detail in Appendix A.2. As shown in Table 1, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases. In addition, we collect corresponding code completion datasets, including Humaneval (Chen et al., 2021), Humaneval-x (Zheng et al., 2023), CodeXGLUE (Lu et al., 2021), MBPP (Austin et al., 2021), and Eval-Plus (Liu et al., 2024). The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios. Currently, most code completion datasets only cover Python, as shown in Table 1. As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
Model Selection. In our experiment, we select three models from two representative Code LLM series, including CodeGemma (Team et al., 2024) and Code-GeeX (Zheng et al., 2023), for evaluation. CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama (Roziere et al., 2023) on code completion tasks. We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales. CodeGeeX series have corresponding applications (cod, 2024) with more than 535k downloads (on VScode (VSC, 2024) Apps shop). We select CodeGeeX2-6b, a version with stable performance. We exclude earlier small-scale models, such as CodeBERT (Feng et al., 2020) and CodeT5 (Wang et al., 2021), because they perform much worse than large-scale models in code completion tasks.
Baseline Selection. We compare our TPIA with GCG-based and manually designed IPI attacks. In particular, the vanilla GCG (Zou et al., 2023) is a representative PI attack and does not meet the requirements of our attack scenarios. Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG). The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work (Greshake et al., 2023). IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments (as shown in Figure 11). Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have some inherent insecurity factors. In particular, we exclude backdoor attacks and adversarial attacks as our baselines because it is difficult to fairly compare these methods with TPIA. Specifically, backdoor attacks require control over the model training process, which is an overly strong assumption for our threat model. Adversarial attacks, on the other hand, cannot achieve target-specific attacks, leading to unreliable threat levels. We provide more explanations in the experiments in Appendix A.3.
Evaluation Metrics. In this paper, we use attack success rate (ASR) to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness. For ASR, we consider an attack successful only when the generated code has equivalent functionality to the target code and appears in the correct position. In other words, the generated code must not only have malicious functionality but also adhere to the code compilation syntax rules. NBR and STF are new metrics introduced for our target-specific attack scenarios. NBR represents the rate of feasible cases. The more cases remaining after excluding the unusable ones with low ASR (below a threshold $r_{1}$), the greater variety of malicious objectives the attack can achieve; STF represents the rate of successful cases and their threat level. The more cases with sufficiently high ASRs (above a threshold $r_{2}$) and larger overall ASR value, the greater the upper limit of the attack’s threat. Specifically, NBR and STF can be formulated as follows:
(7) NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$
STF $=\frac{\sum_{i\in\text{cases where ASR}>r_{2}}\text{ASR}_{i}}{\text{Total \# of cases}},$
where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation. More details about the selection of metrics and values are in Appendix A.2.
Hyperparameters and Computational Facilities. We conduct experiments using the following parameter settings to evaluate our approach. During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100. For forward reasoning enhancement in Section 4.4, the number of tokens reinforced is set to $h=2$, as defined in Equation 6. For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10. Tokenization is performed using Gemma’s tokenizer.
We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization. In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b. These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes.
5.2. Evaluation on Open-source Code LLMs
We first perform the attack evaluation over open-source Code LLMs. This includes three aspects: (1) we assess the attack effectiveness over all evaluated Code LLMs across all threat cases, demonstrating that TPIA is widely effective and more threatening than other attack methods; (2) we evaluate TPIA’s flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
Effectiveness. We hereby evaluate whether TPIA can successfully induce the Code LLMs to generate attacker-specified targets. We control the number of injected tokens for three attack methods in this experiment. For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12. Since IPI-M needs sufficient tokens to be effective, we do not set a fixed length for this attack but limit the number of used tokens under 30 (details of IPI-M are described in Appendix A.2). Table 2 provides the average ASR for four methods. The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code. However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability. On the other hand, the worst-performing CodeGemma-2b is more easily affected by simpler and more direct baseline methods such as C-GCG and IPI-M; Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content. In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M. Therefore, we believe that an increase in the contextual complexity can reduce the ASR of TPIA. In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness. According to our subsequent experiments in Section 5.3 and Section 5.4, increasing the length of TPIA or Top-$k$ value can achieve better attack performance.
An interesting phenomenon is that IPI-M can sometimes enhance the security of the model’s output (producing a lower ASR than ‘No Attack’). We speculate that this is mostly because, in some cases, the generated perturbation contains information on the potential positions where vulnerabilities might occur but does not clearly convey the content of the malicious behavior to Code LLMs. As such, Code LLMs actually strengthen their reasoning about the locations of potential weaknesses and produce more secure outputs. It suggests that (parts of) our TPIA (e.g., those used in IPI-M) could potentially be used for enhancing the security of Code LLMs, by automatically optimizing the unfunctional parts of reference code to improve the security of the generated code. We will further explore it in our future work.
Table 3. The attack evaluation results of NBR and STF against open-source Code LLMs.
Method Metric C/C++ Python Java
  No Attack NBR 0.00 0.22 0.01
STF 0.00 0.11 0.03
C-GCG NBR 0.39 0.50 0.56
STF 0.26 0.47 0.39
IPI-M NBR 0.08 0.25 0.17
STF 0.09 0.19 0.07
TPIA (ours) NBR 0.77 0.92 0.78
STF 0.76 0.87 0.74
Flexibility. The previous experiments show that TPIA has satisfactory effectiveness across all threat cases. However, it is also important to consider the flexibility requirement in practice. In general, it can be divided into two situations. One is when attackers want to achieve diverse malicious objectives, which requires that cases with low ASR be minimized. The other is when attackers want to cause severe damage, which requires that cases with high ASR be maximized. We use NBR to evaluate the attack flexibility for the first situation, and STF for the second situation.
As shown in Table 3, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible. Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat. STF and NBR are not always positively correlated. For example, for TPIA, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score. This indicates that there are more significant and effective attack examples in C/C++. While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results. Besides, whether it is NBR or STF, TPIA significantly outperforms the baselines.
Stealthiness. We hereby evaluate the trade-off between effectiveness and stealthiness. As we discussed in Section 3.3, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals. Arguably, human inspection is currently the only feasible approach to detect TPIA attacks. Therefore, we measure the stealthiness of TPIA by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
As shown in Figure 6, ASR and STF show a clear increase as NIT increases. When NIT is above 12, ASR and STF completely overlap, indicating that ASR exceeds 80% across all cases. Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values. If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy. When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation). This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods, adding just a few tokens can be highly effective.
Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR. This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information. The tokens used by TPIA do not have distinct characteristics in terms of character length but are determined by the attacker’s design. This means that filtering TPIA attacks based on character length is unreliable.
(a) ASR/STF/NIC under Different NIT
(b) NIT/NIC under Different ASR
Figure 6. The study of stealthiness. We plot the above two graphs using NIT and ASR as the x-axis respectively. We sampled 11 different NITs, namely NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
5.3. Evaluation on Commercial Code LLMs
In previous experiments, we have demonstrated that TPIA reaches sufficient effectiveness, flexibility, and stealthiness on open-source models. However, some widely used models are commercial and therefore are closed-source in real-world scenarios. In these cases, attackers need to conduct transfer attacks, i.e., generating perturbations using open-source Code LLMs to attack closed-source ones. We hereby exploit two representative real-world scenarios to evaluate TPIA’s transferability for further verifying its realistic threat.
Attack Scenarios on Real-world Applications. The scenarios with limited access to Code LLM information can be broadly divided into two categories: gray-box and black-box, with the corresponding explanations as follows:
  1. 1)
    Gray-Box Setting. The attacker can infer the open-source model used by the target application through interaction but cannot obtain the exact version of the model. For example, the scale of the model’s parameters, fine-tuning information, and version numbers are unknown to the attacker. We simulate this scenario using CodeGeeX2-6b and the CodeGeeX application. To the best of our knowledge, the CodeGeeX model has at least four different versions, and the CodeGeeX application does not disclose the specific model information it uses. This is a typical gray-box scenario.
  2. 2)
    Black-Box Setting. A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as CodeX). This is precisely the case with GitHub Copilot. We simulate this scenario by using perturbations generated by CodeGeeX2-6b to attack GitHub Copilot.
Table 4. The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
Scenario Language ASR NBR STF
  Gray-box C/C++ 80.4% 0.92 0.75
Python 93.3% 1.00 0.93
Java 71.7% 0.78 0.63
Black-box C/C++ 67.7% 0.70 0.64
Python 90.4% 1.00 0.85
Java 74.4% 0.78 0.65
(a) Perturbation Injection
(b) Target Completion
Figure 7. The example of a successful attack in a real-world black-box scenario (Github Copilot) across files. The perturbation is inserted into the comment of a related file and influence the suggestion of LLM in main file.
Settings. We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total). Other settings are the same as those described in Section 5.2.
Results. As shown in Table 4, in the gray-box scenario, TPIA still demonstrates great attack performance. Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively. Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, TPIA also achieves a satisfying attack performance of 90.4% on Python. The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively. In other words, our TPIA still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent. In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility. Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to real-world commercial applications.
In Figure 7, we present a representative case of ST2 under the black-box scenario. The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function. The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction.
5.4. Ablation Study
In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for TPIA attacks. We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our TPIA.
(a) ASR Study
(b) STF Study
Figure 8. Ablation study of keywords number and Top-$k$. We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords. During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
Effects of Top-$k$. This experiment validates the impact of increasing the Top value on the performance of TPIA. As shown in Figure 8, increasing the Top-$k$ value can significantly improve TPIA’s ASR and STF. Note that, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the vocab size of LLM. However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400). As such, the attacker should assign Top-$k$ value based on their specific needs.
(a) No FRE
(b) FRE
Figure 9. Ablation study of FRE. We compare the differences in ASR and loss function values both with and without using FRE. 1-key1 and 1-key2 represent keeping one of the two keywords, respectively.
Effects of Keyword-based Designing. This experiment evaluates the effects of our optimization module, i.e., keyword-based designing. Specifically, we vary the number of keywords and generate the perturbations. As shown in Figure 8, limiting the number of keywords has a significant impact. The result of 1-keywords perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keywords perturbation. The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%. In contrast, 2-keywords perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95). The above results indicate that using keywords can effectively improve the performance of TPIA.
Effects of Forward Reasoning Enhancement. We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function. Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve. It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion. To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using static instead of optimized keywords. We judge the degree of distortion based on the loss value. The results are illustrated in Figure 9, where we have three key findings. Firstly, under the same conditions, using FRE can lead to a higher ASR. Secondly, without FRE, the attack success rate is still very low ($<20\%$) even when the loss value is lower than that with FRE. Thirdly, FRE has a greater impact on attack success rate than keyword-based design. These results verify the necessity of forward reasoning enhancement in our TPIA method.
6. Discussions
6.1. Resistance to Potential Defenses
In Section 3.3, we have explained why our attack can fundamentally circumvent existing malicious code detection methods. In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack. In general, we can roughly categorize these methods into detection and removal and design following potential defense methods.
Detection: Semantic Analysis. Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code. Since detection methods based on code functionality are ineffective against non-functional perturbations, it is necessary to adopt a semantic analysis approach. This involves examining the differences between the benign code and the suspicious code in the embedding layer of Code LLMs. However, this method may be difficult to implement. We find that code segments injected with perturbations might not have obvious semantic differences from benign ones. Specifically, as shown in Figure 16, we construct classification prompts and extract their embedding layer features on Code LLMs, finding that the two types of code segments are almost completely indistinguishable. We present more details and analyses in Appendix A.3.
Figure 10. The resistance to the removal of chars. We remove characters from the perturbation, selecting 5% for removal each time, and we calculate the average ASR for each programming language across applicable cases. As the influence of the perturbation fades, the ASR results converge to the same level as no attack baseline.
Removal: Massive Deletion and Find-and-Replace. This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code. In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect malicious code segments (as mentioned above). For example, if preserving code readability is a priority, we cannot remove all perturbations, allowing only for a lower degree of modifications. To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations. As shown in Figure 10, if the removal rate of perturbations cannot be ensured to be above 0.5, TPIA still maintains a sufficiently high ASR. On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table 5. Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
Table 5. ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments form.
[…]
Variable Assignment C/C++ 67.5% -8.0%
Python 90.1% -7.8%
Java 78.6% -5.6%
Output Content C/C++ 65.0% -10.5%
Python 83.0% -14.6%
Java 66.5% -9.0%
6.2. Potential Limitations and Future Directions
Although TPIA is effective in attacking the evaluated LLMs and commercial applications, we have to admit that it still has some drawbacks. Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords. We currently select keywords through a simple grid search. How to optimize keywords more efficiently is an important issue. Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent. We believe that readable and benign-looking perturbations may pose a more realistic threat and represent an important direction for future research.
7. Conclusion
In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks. We revealed the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduced TPIA, a new target-specific attack paradigm without controlling the model’s training process. TPIA can circumvent mainstream detection methods (e.g., static analysis and sandboxes) by using non-functional code, instead of functional ones used in software supply chain attacks, to achieve malicious objectives. We evaluated TPIA under three popular programming languages and 31 representative cases. We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation. Our work alerted a new serious security threat in using Code LLMs in practice.
Appendix A Appendix
A.1. Differences with Existing Methods
GCG attack. In the domain of generative models, there exists some adversarial prompt injection attack methods, such as universal trigger (Wallace et al., 2019) and well-known GCG (Zou et al., 2023). Taking GCG as an example, the adversarial goal of the attacker is to force LLMs to generate affirmative responses beginning with “Sure, here’s…,” thereby disrupting security alignment (Ji et al., 2024). Unlike GCG, TPIA emphasizes the feasible usage of perturbations and performance during indirect attacks. First, TPIA transfers different malicious objectives to perturbations to induce specific code generation, while GCG aims to generate the most powerful perturbation to break the safety guard. Second, TPIA considers a scenario where the users are victims and uses keyword-based perturbation to save the number of perturbation tokens, while GCG assumes that the users of LLMs are malicious attackers and allow unlimited tokens. Third, TPIA proposes a code-oriented task construction framework and forward reasoning enhancement technique to ensure effectiveness and robustness in uncertain contexts, while GCG has a controllable context and fixes the perturbation at the end of the prompt. What is more, straightforward application of adversarial perturbation methods like GCG on code completion attack has a terrible performance. As shown in Figure 9, 10-adv attacks with “0-keyword no FRE” refers to the 0% ASR.
Codebreaker and SVEN. In the domain of code, the newly proposed backdoor attack method, Codebreaker (Yan et al., 2024), also leverages the feedback of LLMs for optimization. However, unlike TPIA, Codebreaker emphasizes using the coding capabilities of LLMs to reduce the probability of the malicious payload code being detected, thereby increasing the success rate of backdoor attacks, without removing the reliance on the poisoning process. A method called SVEN (He and Vechev, 2023) involves purposefully generating high-risk code. However, this method requires Prefix-tuning on Code LLMs, adding new parameters to the original model parameters, making it unsuitable as an attack method. Additionally, when generating high-risk code, SVEN cannot control the content of risk code, nor can it achieve target-specific attacks.
A.2. Additional Details
Figure 11. Greshake et al. (Greshake et al., 2023) use a large amount of disruptive information and explicit malicious code snippets in the comments, which has a significant impact on the feasibility of the attack.
Baseline Details. IPI-M and C-GCG are important baselines in our effectiveness experiment. IPI-M uses injection instructions in natural language form to interfere with the model’s reasoning results, commanding the model to generate insecure code. The original IPI-M places a large amount of disruptive commentary along with the malicious code to be inserted in the comment section, as shown in Figure 11. We argue that this method has very low feasibility. Despite the attack path of IPI does not need to worry about being detected by malicious code detection methods, there is a risk of being discovered and removed once the victim becomes aware and actively pays attention to IPI attack, especially perturbations with poor visual stealthiness. Therefore, we attempt to simplify IPI-M into concise and clear template-based injection sentences, as shown in the table. C-GCG mainly follows the GCG method, using an adversarial prefix to achieve malicious purposes. At the same time, C-GCG does not adopt the two optimization methods from Section 4.4 and Section 4.5. C-GCG performs much worse than TPIA in the main experiments.
Value and Metric Selection. We select the values of $r_{1}$ and $r_{2}$ based on the attack’s performance in the experiment. Specifically, in our experiments, 20% ASR is smaller than the ASR of no attack baseline in some situations, and 80% ASR outperforms the best performance of all the baselines. In addition, we find that only a few subcases (determined by the model, dataset, attack method, and case) have a success rate between 20% and 80%, accounting for 7.5%. Therefore, setting 20% and 80% as threshold of bad case and outstanding case respectively is appropriate.
As for metrics, we have already explained why metrics such as NBR and STF are used in the main experiments. Here we further explain some details related to them. First, NBR does not take the specific ASR value into account, as this metric is used to measure the diversity of attacks, rather than ensuring stable success rates. Second, STF targets the situation where attackers aim to cause as much damage as possible, and it varies with differences in ASR. In other words, even a difference between 100% ASR and 95% ASR can have a real impact on the overall threat level. Therefore, the calculation of STF considers the specific ASR value. In addition, we do not use common metrics for code generation tasks, such as BLEU or CodeBLEU (Ren et al., 2020), because they do not directly reflect the effectiveness of the attacks. Overall, we attempt to confirm the realistic meaning of each evaluation metric from different perspectives.
Case Study. Here we show the details of our malicious objectives and threat cases. As shown in Table 6, we design 13 malicious objectives and 31 threat cases in 3 programming languages based on CWE top-10 KEV, exisiting research and famous destructive instruction. In programming process, minor changes in the code can cause significant functional alterations. We roughly divide the causes of vulnerabilities in all malicious objectives into three categories: skipping critical process (CWE20, 78, 502, 918, 22, 306), wrong memory usage (CWE416, 122, 787, 843), and outputting dangerous snippets (ST0, 1, 2). All these vulnerabilities result in severe consequences.
Table 6. Description of Threat Cases.
Case Description From
CWE416 Use after free. CWE 1st KEV
CWE122 Heap-based buffer overflow. CWE 2nd KEV
CWE787 Out-of-bounds write. CWE 3rd KEV
CWE20 Improper Input Validation CWE 4th KEV
CWE78 ‘OS Command Injection’ CWE 5th KEV
CWE502 Deserialization of untrusted data. CWE 6th KEV
CWE918 Server-side request forgery. CWE 7th KEV
CWE843 Access of resource using incompatible type. CWE 8th KEV
CWE22 ‘Path Traversal’ CWE 9th KEV
CWE306 Missing authentication for critical Function. CWE 10th KEV
ST0 Use ECB mode to encrpty. Schuster et al.(Schuster et al., 2021)
ST1 Use SSL3 protocol Schuster et al.(Schuster et al., 2021)
ST2 Executing “rm -rf / Famous destructive instruction
(a) Skipping critical process - CWE20
(b) Wrong memory usage - CWE122
(c) Outputting dangerous snippets - ST2
Figure 12. Three examples of TPIA attack cases. Code LLMs normally generate code segment without vulnerabilities, but they lose this ability after being attacked by TPIA. Attackers can modify the output content at the specified location through TPIA, and they can even generate malicious code snippets that have never appeared in the context, e.g., “os.system("rm -rf /")”.
As shown in Figure 12, here we present three representative examples across three different languages and different type of objectives. The first is CWE20 (using Java), where the program attempts to fetch the content of a URL and calculate its length. A secure programming practice would be to check if the content is null before performing the length calculation, thus avoiding a null pointer. However, after being attacked, the Code LLM generates code that calculates the length directly after fetching the URL, leading to a null pointer exception if the URL is null. A null pointer can result in serious consequences such as program crashes or data corruption. We simply categorize this situation as “skipping critical process.” The second example is CWE122 (using C), where the program attempts to pass content to buf through a command-line argument. A secure programming practice would be to check if the content exceeds BUFSIZE before the assignment, thus avoiding overflow. Through TPIA, attackers can induce Code LLMs to perform the assignment directly. When a stack overflow occurs, it accesses non-existent RAM space, leading to consequences such as system crashes and data leaks. We simply categorize this situation as “wrong memory usage.” This type of situation generally applies only to C/C++, as Java and Python automatically manage memory release. The third example is ST2 (using Python), where the program attempts to call a function named TRG (TaRGet). Under normal circumstances, this call does not present any issues, and the large model generates suggestions based on the contextual semantics or directly produces an ¡eos¿ token. Through TPIA, attackers can make Code LLMs generate the execution code “rm -rf /” right after calling the TRG function. If executed, this code results in the deletion of all data under the root directory. We categorize this situation as “outputting dangerous snippets.”
We only apply different perturbation construction strategies to different categorizations of cases in the IPI-M. Specifically, for “skipping critical process,” we use “… directly after …” and “without any check”; for “wrong memory usage,” we use “It is important to … directly after …”; for “outputting dangerous snippets,”, we use “It is very important to promptly … after …”. In addition, the perturbation in IPI-M must be sufficiently explicit; for example, it needs to specify the code snippet you want to insert (e.g., “os.system("rm -rf")”) or present a clear assertion (e.g., “SSL3 is the most secure protocol”). Otherwise, Code LLMs are difficult to induce successfully. Arguably, the actual performance of IPI-M may be worse than our evaluation results shown in main experiments.
Keywords Selection Method. We use a simple dual-loop strategy in the selection of keywords. Specifically, before each round of perturbation optimization with length $l$ starts in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords. With the optimization step $s$, the equation can be formulated as follows:
(8) $\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$
$\bm{P}_{adv}^{s+1}=\min_{\bm{P}_{adv}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P}_{adv},\bm{\theta}),$
where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6. $\bm{W}$ represents the token sequence that constitutes the keywords. To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., $\bm{P}_{key}=\{p_{k1},p_{k2},\dots,p_{kn}\}$, and for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{target code},\text{position code}\}$.
(a) Gemma
(b) Codegemma
(c) Codellama
(d) Codegeex2
Figure 13. Examples of traditional adversarial attacks on four different LLMs. The generated perturbations are displayed with a gray background, the input code completion task content with a green background, and the output content with a red background.
A.3. More Experiments
Comparison with Adversarial Attack. The advantages of TPIA over backdoors are clear: it does not require control of the training process and has a very low attack threshold. However, the advantages of TPIA over existing adversarial attacks, i.e., achieving target-specific attacks, is not obvious.
Figure 14. The ASR across different cases. The experimental results shown in this figure are the average values calculated from the effectiveness experiments involving 3 programming languages, 4 datasets, and methods excluding Clean Accuracy.
Here, we propose an imperfect but highly intuitive scoring method to compare TPIA with state-of-the-art adversarial attack method (Jha and Reddy, 2023). We divide the output results into three levels: (1) generating code with normal functionality, with a harmfulness score of 0 (level 0); (2) generating code that cannot compile or has functional errors, with a harmfulness score of 0.5 (level 1); (3) generating code with stealthy risk, with a harmfulness score of 1 (level 2). In terms of the resulting effects, the second level will lead to a decrease in the usability of Code LLM, and the third level not only affects the usability but also causes unpredictable financial losses to the users of the Code LLMs. We evaluate the performance on three cases: S0, S1, and S2, using four models: Gemma-7b, CodeGemma-7b, CodeLlama-7b, and CodeGeeX-6b (Due to the significant differences in resistance to traditional adversarial attacks across different models, we use more types of LLMs than in the main experiments, including a general-purpose LLM).
Figure 15. A comparison of the harmfulness between CodeAttack (CA) and TPIA. The final scores were normalized to the 0-100. We use light colors (CA lv1 and TPIA lv1) to represent ratings from level 1 harmfulness and dark colors (CA lv2 and TPIA lv2) to represent ratings from level 2 harmfulness.
Figure 16. The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b. The two types are almost indistinguishable.
As shown in Figure 15, TPIA significantly outperforms CodeAttack in scores. This is because TPIA consistently achieves level 2 harmfulness, whereas CodeAttack tends to reach level 1. In fact, the final result of CodeAttack is not invalid: the quality of output is greatly impacted, as shown in Figure 13. However, the output code usually cannot be correctly compiled. In other word, this type of attack is unlikely to threaten the user’s experience, although it successfully reduces the usability of Code LLM and has a negative impact on its commercial applications. In contrast, TPIA also impairs the usability of the victim Code LLMs, and can cause terrible subsequences for victim users. In summary, TPIA is capable of conducting more harmful attacks with a higher success rate, making it more suitable for LLM-based code completion scenarios compared to traditional adversarial attack methods.
Effectiveness across Different Cases. In the main experiments, we have validated the outstanding attack performance of TPIA from three perspectives: model, dataset, and programming language. Here, we supplement the attack details for different attacks on a single case. Specifically, we use ASR to evaluate the performance of TPIA and the baselines across different cases. We present the specific average ASR for each case. As shown in Figure 14, TPIA achieves a certain level of effectiveness across all cases, with an overall average ASR of 83.4%. However, C-GCG and IPI-M both have cases where they are completely ineffective, especially for IPI-M. This demonstrates that directly using traditional GCG methods or manually designing perturbations is not effective in some scenarios. We believe this involves three reasons: (1) The baseline methods are not fully suitable for contexts with structured features, which is a fundamental characteristic of the code completion task; (2) The specific malicious objective include complex semantics such as target code, position code, and “disregarding security attributes,” which are sometimes not fully captured by the perturbations generated by the baseline methods; (3) The impact of keyword-based designing and FRE; our ablation study proves that these two modules play a critical role in TPIA’s effectiveness.
According to the results, the most challenging to achieve is the straightforward case ST2. This case requires generating an obviously malicious command in a completely unrelated context. Another difficult case to implement is CWE22 “Path Traversal,” which involves using unconventional paths like “../” to access files that are normally inaccessible. This case is relatively easier to achieve in Python, where the Code LLMs tend to suggest directly reading the path. However, it is more challenging to implement in Java and C/C++, as Code LLMs often only recommend checking the path rather than directly reading the file. CWE843 involves manipulating union members in a struct in C/C++. This represents a more complex malicious objective, where we attempt to manipulate one value within the union using an untrusted value. This is an unconventional operation that is difficult to describe using natural language, making it challenging for IPI-M to successfully attack. In contrast, methods like TPIA and C-GCG, which convey malicious instructions through token selection, are able to achieve success. Arguably, TPIA has a clear advantage in attack performance for all the cases we have collected.
Semantic Analysis on Payload Code. In this experiment, we attempt to detect malicious payload code by semantic analysis. We observe the classification features of code segment in the embedding layer of CodeGemma-2b. Specifically, we add the classification prompt keyword [CLS] to 820 original code snippets and 820 malicious payload code snippets totally. and then input them into the model to analysis their semantic features. We examine the t-SNE (Van der Maaten and Hinton, 2008) distribution of the classification keyword [CLS] in the embedding layer, which is a widely used method for classification tasks (Abdelnabi et al., 2024). As shown in Figure 16, the features of the two types of text are almost completely intertwined, making it nearly impossible to distinguish between the injected code data and the original code data. In addition to observing the characteristics of layer 17 shown in Figure 16, we also examine the data features of layers 1 through 16. There is no significant difference observed in these layers compared to layer 17. There are several reasons for this situation. First, code data, including comments, tend to have lower fluency and coherence with high token complexity, which poses a significant challenge for semantic analysis tasks. Second, the perturbations used by TPIA occupy an extremely small number of tokens, exerting only a minor influence on the overall semantics of the injected data, making them hard to detect. Third, the semantic impact of TPIA differs from traditional injection attacks in that it does not require an extremely high intensity of information to counteract the model’s secure alignment; instead, it guides the model to generate vulnerabilities that could occur naturally, thus making detection through semantic analysis difficult. Our experiment results indicate that simple semantic analysis methods are not effective in defending against TPIA. This indicates that detecting code snippets containing malicious perturbations generated by TPIA through semantic analysis is not an effective approach.

### v5

@@
-TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented Large Language Models
-Abstract.
+TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs
+Abstract
 Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming.
-With these tools, developers can easily generate the desired complete functional code based on incomplete code snippets and natural language prompts.
 Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks.
@@
 The latter crafts malicious adversarial input codes to reduce the quality of the generated codes.
-However, both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes.
-This paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA).
+In this paper, we reveal that both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes.
+To alleviate these problems, this paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA).
 TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base).
@@
 We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
-Our work alerts a new practical threat of using Code LLMs.
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
-1. Introduction
+I Introduction
 Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation.
-In particular, code-oriented large language models (Code LLMs), e.g., CodeX (Wojciech Zaremba, 2021), CodeGemma (Team et al., 2024), and CodeGeex (Zheng et al., 2023), have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus (Zhang et al., 2023; Li et al., 2023b; Ding et al., 2024) and the fundamental paradigm in commercial practice (Friedman, 2021; cod, 2024)..
+In particular, code-oriented large language models (Code LLMs), e.g., CodeX [53], CodeGemma [49], and CodeGeex [61], have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus [60, 34] and the fundamental paradigm in commercial practice [19, 2]..
 They can efficiently generate complete functional code based on incomplete code snippets and text prompts.
-Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot (Friedman, 2021) developed jointly by Github (Git, 2024) and OpenAI (Ope, 2024).
+Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot [19] developed jointly by Github [3] and OpenAI [4].
 These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
@@
 In general, existing attacks can be divided into two main categories.
-(1) Backdoor attacks (Schuster et al., 2021; Li et al., 2023a; Yang et al., 2024).
-They intend to compromise the models by poisoning the training data (Schuster et al., 2021; Yang et al., 2024) or model parameters (Li et al., 2023a), such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns.
-They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications (Li et al., 2023a), or implementation of less secure encryption rules or protocols (Schuster et al., 2021).
-(2) Adversarial attacks (Zhou et al., 2022; Bielik and Vechev, 2020; Jha and Reddy, 2023).
+(1) Backdoor attacks [46, 29, 57].
+They intend to compromise the models by poisoning the training data [46, 57] or model parameters [29], such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns.
+They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications [29], or implementation of less secure encryption rules or protocols [46].
+(2) Adversarial attacks [62, 12, 24].
 They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs.
-For example, Jha et al.(Jha and Reddy, 2023) slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score (Ren et al., 2020).
+For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs.
 However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent.
@@
 Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult.
-Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks (Bielik and Vechev, 2020; Zhou et al., 2022; Jha and Reddy, 2023) can hardly achieve precise malicious objectives, although they do not require intervention in the training process.
+Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks [12, 62, 24] can hardly achieve precise malicious objectives, although they do not require intervention in the training process.
 This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input.
@@
 Unfortunately, the answer to the above question is negative.
-In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger (Wallace et al., 2019) and Greedy Coordinate Gradient (GCG) (Zou et al., 2023).
-Please find more details in Section 2 and Appendix A.1..
+In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger [50] and Greedy Coordinate Gradient (GCG) [63].
+Please find more details in Section II..
 TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks.
@@
 Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice.
-For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs (Chen et al., 2024) via retrieval-augmented generation (Lewis et al., 2020).
-In particular, different from supply chain attacks (Ohm et al., 2020) that can be easily detected by static analysis (Li et al., 2018b) or sandboxing (Alhamdan and Staicu, 2023) since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
-Figure 1.
-The attack paragdim of our TPIA.
+For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27].
+In particular, different from supply chain attacks [41] that can be easily detected by static analysis [33] or sandboxing [9] since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
+Figure 1: The attack paragdim of our TPIA.
 A concise non-functional attacker-generated perturbation is injected into one of the project’s files.
@@
 In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective.
-Specifically, we obtain the base Code LLM of the integrated application and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code.
+Specifically, we obtain the base Code LLM and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code.
 This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness.
-(2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack.
-We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) (Top, 2024) across 3 programming languages.
+(2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the number of tokens required for a successful attack.
+We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages.
 There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++.
@@
 The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness.
-In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX (cod, 2024) (with more than 535k installations in the VScode (VSC, 2024) extension shop) and Github Copilot (Friedman, 2021) (developed by Github and OpenAI (Ope, 2024)).
+In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 535k installations in the VScode [6] extension shop) and Github Copilot [19] (developed by Github and OpenAI [4]).
 These applications correspond to the gray-box and black-box attack scenarios, respectively.
@@
   • 3)
-    We conduct extensive experiments to evaluate our TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 widely used commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
-2. Background and Related Work
-2.1. Code LLMs for Code Completion
-Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation (Sun et al., 2022; Roziere et al., 2020), and code summarization (McBurney and McMillan, 2016; Ahmad et al., 2020).
-Many models have been developed to achieve these tasks, including CodeGen (Nijkamp et al., 2022), CodeBERT (Feng et al., 2020), and CodeT5 (Wang et al., 2021).
+    We conduct extensive experiments to evaluate TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
+II Background and Related Work
+II-A Code LLMs for Code Completion
+Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation [48, 45], and code summarization [40, 8].
 The emergence of LLMs provides new alternatives to the coding tasks.
-By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex (Wojciech Zaremba, 2021), CodeLlama (Roziere et al., 2023), CodeGemma (Team et al., 2024) and CodeGeex (Zheng et al., 2023).
-These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast array of functionally complex code (Xia et al., 2023).
-In this paper, we focus on code completion (Li et al., 2023b; Ding et al., 2024; Zhang et al., 2023).
-This is a paradigmatic task that involves generating new code from incomplete input (with natural language) using Code LLMs, which have been integrated into commercial applications (Friedman, 2021; cod, 2024).
+By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [53], CodeLlama [44], CodeGemma [49] and CodeGeex [61].
+These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast of functionally complex code [54].
+In this paper, we focus on code completion [34, 16, 60].
+This is a paradigmatic task that involves generating new code from incomplete input using Code LLMs, which have been integrated into commercial applications [19, 2].
 However, this development paradigm may encounter some potential threats.
@@
 On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs.
-These threats have already occurred in the real world (Cri, 2025), highlighting the necessity of its security research.
-2.2. Backdoor and Adversarial Attacks
-Backdoor attacks target the training process of DNNs (Liu et al., 2020; Li et al., 2022; Zhang et al., 2024).
-These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present.
-In the code domain, there are generally two strategies to achieve this, including dataset poisoning (Schuster et al., 2021; Yang et al., 2024) and model poisoning (Li et al., 2023a).
-The former requires the attacker to modify the code dataset used for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM used for application development.
-However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned research institutes or companies).
+These threats have already occurred in the real world [7], highlighting the necessity of its security research.
+II-B Backdoor and Adversarial Attacks
+Backdoor attacks target the training process of DNNs [20, 31, 58].
+These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [26, 30, 21].
+In the code domain, there are generally two strategies to achieve this, including dataset poisoning [46, 57] and model poisoning [29].
+The former requires the attacker to modify the code dataset for model training, while the latter requires the attacker to modify the parameters of the pre-trained Code LLM for application development.
+However, in real-world scenarios, developers of Code LLMs or integrated applications usually obtain the datasets or pretrained model from trusted entities (e.g., renowned institutes or companies).
 This mitigates the backdoor threats from the source.
-Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult.
-Arguably, backdoors cannot be implanted into a Code LLM without controlling the model training process.
-Different from backdoor attacks, adversarial attacks primarily target the model’s inference process (Li et al., 2018a; Doan et al., 2021; Wang et al., 2022).
-Adversarial attacks induce the vulnerabilities of models by introducing well-crafted perturbations to the input data, leading the model to make incorrect output.
-In the code domain, existing adversarial attacks (Bielik and Vechev, 2020; Zhou et al., 2022; Jha and Reddy, 2023) aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs.
-For example, Jha et al.(Jha and Reddy, 2023) slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score (Ren et al., 2020).
+Recently, a few works also exploited backdoor attacks for positive purposes [32, 55, 52], which are out of our scope.
+Adversarial attacks target the model’s inference process [11, 43, 25].
+Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [59, 28, 47].
+In the code domain, existing adversarial attacks [12, 62, 24] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs.
+For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [42].
 However, existing adversarial attacks are difficult to fulfill specific malicious purposes since they can hardly achieve precise malicious objectives, although they do not require manipulation in the training process.
 This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, inevitably weakening Code LLMs’ understanding of the input.
-Accordingly, adversarial attacks are usually difficult to pose a serious threat to Code LLMs in practice.
-2.3. Prompt Injection and Indirect Prompt Injection Attacks
-Prompt injection (PI) attacks, are a special form of attack targeting LLMs (DAN, 2023; Liu et al., 2023b; Zou et al., 2023).
+Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
+II-C Prompt Injection and Indirect Prompt Injection Attacks
+Prompt injection (PI) attacks, are a special form of attack targeting LLMs [15, 36, 63].
 A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb.
-A general way to carry out these attacks is by submitting a carefully designed prompt query to the LLM.
-In the code domain, existing research (Jiang et al., 2023; Liu et al., 2023a; Fang et al., 2024) focuses on code-oriented attack objectives, such as using LLMs to create malware.
+In the code domain, existing research [39, 37, 17] focuses on code-oriented attack objectives, such as using LLMs to create malware.
 In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
-Indirect prompt injection (IPI) attack (Greshake et al., 2023), on the other hand, considers a different attack scenario than PI.
-The attacker attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites or providing false information.
+Indirect prompt injection (IPI) attack attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites [22].
 This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses).
-In the code domain, the only existing work (Greshake et al., 2023) makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it needs to insert a large amount of readable comments containing the malicious intent, which can be easily noticed by programmers, as shown in Figure 11.
+In the code domain, the only existing work [22] makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it does not consider stealthiness and inserts a large amount of readable comments containing the malicious intent.
 Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
-3. Problem Statement
-3.1. Code Completion Task
+III Problem Statement
+III-A Code Completion Task
 Code completion is an essential task for Code LLMs.
@@
 We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
-3.2. Target-specific Attack
+III-B Target-specific Attack
 In the code domain, target-specific attack refers to the scenario where the attacker makes the Code LLM generate unsafe code for specific targets.
@@
 To conclude, for a successful target-specific attack, the model output received by the victim $\bm{Y_{v}}$ will be:
-(1) $\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\}.\end{split}$
-3.3. Threat Model
+$\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\},\end{split}$ (1)
+where $\oplus$ means contacting the two token sequences.
+Figure 2: The overall pipeline of our TPIA attack.
+It is executed with two attack steps.
+(1) Objective Transfer.
+The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective.
+(2) Perturbation Injection.
+The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub [3] and StackOverflow [5]) and introduced to the victim’s code context via various paths, such as direct copy and installing packages.
+When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
+III-C Threat Model
 Attack Scenarios.
-We consider common scenarios where users introduce external code by copying or installing packages.
-The attacker generates a concise non-functional perturbation containing malicious instructions and inserts it into the victim’s code context via the external code.
-When the user requests the Code LLM to perform code completion based on code context, the perturbation will induce the model to generate an unsafe output at the specified location.
-In particular, different from supply chain attacks (Ohm et al., 2020) that can be easily detected by existing main-stream detection methods (e.g., static analysis (Li et al., 2018b)) since they directly insert malicious functional code, to implant malicious behaviors, our attack scenario is more threatening because the attacker can circumvent detection methods by introducing only non-functional malicious code.
-Besides introducing code to the victim’s local project, attackers can upload their malicious payload codes to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs (Chen et al., 2024) via retrieval-augmented generation (Lewis et al., 2020).
-Code in the knowledge base can also appear in the context of the Code LLM’s prompts.
+We consider common scenarios where users introduce external code by copying or installing pack- ages [22, 41].
+The attacker generates a non-functional perturbation to circumvent detection methods (e.g., static analysis [33] and sandbox [9]) and inserts it into victim’s code context via the external code.
+When the user requests the Code LLM to perform code completion, the perturbation will induce the model to generate an unsafe output at the specified location.
+Besides introducing code to the victim’s local project, attackers can also upload their codes containing perturbation to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27].
+Perturbation in the knowledge base can also appear in the context of the Code LLM’s prompts.
 Attacker’s Capabilities.
-We assume that the attacker can obtain the parameters of Code LLM to generate malicious perturbations.
-Arguably, when developers build integrated applications using open-source models (such as CodeGemma (Team et al., 2024), CodeLlama (Roziere et al., 2023), or CodeGeeX (Zheng et al., 2023)), obtaining the corresponding model parameters is feasible.
-In this scenario, TPIA can achieve optimal performance.
-In this paper, we also consider gray-box (i.e., with unknown model version and system prompts) and black-box (i.e., with closed-source commercial models) scenarios, based on two widely-used commercial Code LLM-integrated applications in Section 5.3.
-We conduct successful attacks against these applications via perturbations generated with a local open-source Code LLM, confirming the practical threat of our TPIA in the real world.
-On the other hand, the attacker needs to submit the payload code into the code repository and prevent it from being detected.
-Since our threat model uses non-functional code to achieve objectives, the payload code does not need to be designed with malicious functionality.
-In other words, the attacker can freely design the functional part of the payload code to ensure successful submission without worrying about detection by methods based on malicious functionality.
-In contrast, the perturbation part of the payload needs to be stealthy enough to avoid detection by victim users.
+We assume that the attacker can obtain the parameters of an assistive Code LLM to generate malicious perturbations.
+This assumption is consistent with many adversarial methods [63, 50, 24].
+Specifically, the attacker ideally obtains the open-source model corresponding to the target application, or uses the most effective assistive model in gray-box and black-box scenarios.
+On the other hand, perturbations can be detected and removed.
+Therefore, the attacker needs to employ non-functional perturbations rather than the functional code commonly used in supply chain attacks [41], since the latter can be easily detected by existing methods (e.g., static analysis [33] and sandboxing [9]).
 Attacker’s Goals.
 The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code).
-This tuple follows the definition of target-specific attacks in Section 3.2.
+This tuple follows the definition of target-specific attacks in Section III-B.
 Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$.
@@
 It can be formulated as below:
-(2) $\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$
-Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section 3.2 and the generated perturbation $\bm{P}$ should satisfy Equation (1).
+$\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$ (2)
+Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section III-B and the generated perturbation $\bm{P}$ should satisfy Equation (1).
 In particular, the attacker can achieve these goals by inserting code snippets into specified positions, without needing to oppose the Code LLMs’ safety alignment.
 For example, when setting the target position to DES.new( and the target code snippet to key,DES.MODE_ECB), the generated perturbation can induce Code LLMs to use the easily cracked ECB mode for encryption; When setting the target position to TRG() (a shorthand for Target which can be replaced with any function name), the generated perturbation can induce Code LLMs to insert designed code snippet into the TRG function or after using it.
-Figure 2.
-The overall pipeline of our TPIA attack.
-It is executed with two attack steps.
-(1) Objective Transfer.
-The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective.
-(2) Perturbation Injection.
-The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub (Git, 2024) and StackOverflow (Sta, 2024)) and introduced to the victim’s code context via various paths, such as direct copy and installing packages.
-When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
-4. Methodology
-4.1. Overview
-We introduce our target-specific prompt injection attack (TPIA).
+IV Methodology
+IV-A Overview
+We hereby briefly introduce our target-specific prompt injection attack (TPIA).
 In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2.
-Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section 4.2).
-In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section 4.3).
-The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section 4.4); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section 4.5).
+Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section IV-B).
+In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section IV-C).
+The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section IV-D); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section IV-E).
 We describe their technical details in the following parts.
-4.2. Task Construction
+IV-B Task Construction
 In the first stage, TPIA constructs specific code completion tasks based on the malicious objectives.
@@
 The response of the Code LLMs should be the target position.
-As stated in Section 3, the target code should be inserted after the target position to achieve the malicious objective.
+As stated in Section III, the target code should be inserted after the target position to achieve the malicious objective.
 Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
@@
 In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”.
-Finally, TPIA constructs the code unrelated to the task, such as “generate_random_string”, and uses it as the noise-like context code.
+Finally, TPIA constructs the code unrelated to the task and uses it as the noise-like context code.
 Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
-Figure 3.
-The example of task construction that induces Code LLM to use encryption mode with low security.
+Figure 3: The example of task construction that induces Code LLM to use encryption mode with low security.
 In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
-4.3. Malicious Perturbation Generation
+IV-C Malicious Perturbation Generation
 In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code.
-Utilizing the code completion task constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer.
+Utilizing the task code constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer.
 Through a greedy search approach, the optimal perturbation is derived.
@@
 Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation.
-Considering that our enhancement method adds static tokens to the perturbation (Section 4.5), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$.
-After determining the input and output of the code completion task, TPIA performs continuous iterations to optimize the perturbation.
-Specifically, since the attack is ultimately executed by the perturbation embedded in the external source code, in each iteration, TPIA only replaces the adversarial part of the perturbation token by token (Line 4).
+Considering that our enhancement method adds static tokens to the perturbation (Section IV-E), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$.
+After determining the input and output of the code completion task, TPIA starts to optimize the $x_{\text{adv}_{1:l}}$ token by token in each iter (Line 4).
 We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
-(3) $\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$
+$\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (3)
 where $\bm{T_{s}}$ denotes the task code.
@@
 Besides, the optimization objective for the $i$-th adversarial token $x_{\text{adv}_{i}}$ can be denoted as:
-(4) $\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$
+$\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$ (4)
 To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer.
-Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
-(5) $e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$
+Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set of the Code LLM $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
+$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$ (5)
 where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss.
 TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7).
-This has been proven to be highly effective in existing NLP tasks (Zou et al., 2023; Wallace et al., 2019).
+This has been proven to be highly effective in existing NLP tasks [63, 50].
 Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9).
@@
 Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
-In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger (Wallace et al., 2019) and GCG (Zou et al., 2023).
+In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger [50] and GCG [63].
 We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection).
@@
 To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA.
-More details about the comparison between existing methods and TPIA are discussed in Appendix A.1.
-Figure 4.
-Loss function distortion during the forward reasoning process.
+Figure 4: Loss function distortion during the forward reasoning process.
 When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence.
-However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective.
-4.4. Forward Reasoning Enhancement
+However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective in practical scenarios.
+IV-D Forward Reasoning Enhancement
 This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness.
-Specifically, the calculation of the loss function in Malicious Perturbation Generation will encounter the distortion issue during the actual reasoning process of the Code LLMs.
 As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations.
 For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context.
-In our experiment, a malicious perturbation might have a much smaller loss value compared to a successful manually-crafted perturbation, yet the former fails to produce the desired output.
 This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions.
@@
 Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error.
-Its formulation is shown as follows:
-(6) $\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$
+Its specific formulation is shown as follows:
+$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (6)
 $\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
 where $\mathcal{L}_{e}$ represents the additional reasoning-based enhancement applied to the original loss function, and $h$ is a hyperparameter set to either 1 or 2, which significantly improve the performance of attack in practice.
-Figure 5.
-Examples of the keyword-based perturbation and other perturbations.
+Figure 5: Examples of the keyword-based perturbation and other perturbations.
 The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage.
 Our keyword-based perturbation can balance both effectiveness and stealthiness.
-4.5. Keyword-based Perturbation Design
+IV-E Keyword-based Perturbation Design
 This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack.
-Specifically, since the content of the perturbation and the payload code is not logically consistent, the attacker must enhance perturbation’s stealthiness as much as possible to avoid human detection.
 From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack.
 Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent.
-However, although we can use some optimization methods to obtain the optimal perturbation, it is still difficult to use illegible tokens to express complex instructions.
-A typical failure case is that even after the optimization, the attack objective cannot be completely transferred into the perturbation, resulting in attack invalidation.
+However, accurately conveying instructions using illegible tokens remains a significant challenge.
 A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues.
@@
 As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds.
-Additionally, the complex tokens generated after adversarial computation often occupy more vision space.
 To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design.
-This method allows the malicious perturbation to consume fewer tokens to convey malicious instruction, enhancing the attack stealthiness.
 Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens.
-Specifically, we need to use non-sensitive information that does not directly contain malicious features, which we call keywords, as static tokens for perturbation.
-For instance, as shown in Figure 5, part of the target function name (TR) is the key information, which is not directly related to the malicious objective (data deletion).
-The tokens of keywords are selected from the position code and target code after each perturbation optimization process, and the content is ensured to be non-malicious by limiting the number and position of selected tokens.
-The details of this method is discussed in Appendix A.2.
-By initializing the perturbation as “keywords<AT><AT><AT><AT>”, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
-5. Evaluation
-In this section, we conduct an extensive evaluation of TPIA, demonstrating its threat to Code LLMs from the following perspectives:
+Specifically, before each round of perturbation optimization in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords.
+With the optimization step $s$, the equation can be formulated as follows:
+$\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$ (7)
+$\bm{P}_{adv}^{s+1}=\min_{\bm{P}_{adv}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P}_{adv},\bm{\theta}),$
+where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6.
+To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{\emph{target code}},\text{\emph{position code}}\}$.
+As shown in Figure 5, part of the target function name (TR) is the keywords, which is not directly related to the malicious objective (data deletion).
+By using a small number of statistical tokens, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
+V Experiments
+In this section, we conduct an extensive evaluation of our TPIA, demonstrating its realistic threat to Code LLMs from the following three perspectives:
   1. 1)
-    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section 5.2)
+    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section V-B)
   2. 2)
-    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section 5.3)
+    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section V-C)
   3. 3)
-    TPIA’s modules are all effective in improving attack performance. (Section 5.4)
-5.1. Experiment Settings
-Table 1.
-The summary of objectives and datasets.
+    TPIA’s modules are all effective in improving attack performance. (Section V-D)
+V-A Main Settings
+TABLE I: The summary of objectives and datasets.
 We list the programming languages that each objective/dataset is applicable: C/C++ (C), Python (P), Java (J).
@@
 Data Count: 14273 (C/P/J): 2132/7965/4176
-Table 2.
-The ASR across different Code LLMs.
+TABLE II: The ASR across different Code LLMs.
 The experiment targets four code completion datasets and tests the success rates of various attack methods using three Code LLMs.
-As shown in Table 1, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets.
+As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets.
 Best attack results are marked in bold.
@@
   CodeGemma-2b No Attack 1.9% 17.3% 9.4% 22.0% 4.9% 12.6% 17.0% 12.2%
-C-GCG (Zou et al., 2023) 40.6% 50.2% 32.1% 36.5% 49.4% 59.0% 44.2% 44.6%
-IPI-M (Greshake et al., 2023) 3.4% 30.8% 6.4% 33.3% 3.2% 19.7% 31.0% 18.3%
+C-GCG [63] 40.6% 50.2% 32.1% 36.5% 49.4% 59.0% 44.2% 44.6%
+IPI-M [22] 3.4% 30.8% 6.4% 33.3% 3.2% 19.7% 31.0% 18.3%
 TPIA (ours) 75.5% 93.4% 84.2% 74.4% 67.3% 94.1% 85.4% 82.0%
 CodeGemma-7b No Attack 0.0% 12.2% 0.0% 18.9% 0.0% 15.6% 12.1% 8.4%
-C-GCG (Zou et al., 2023) 43.5% 53.8% 55.4% 33.1% 20.4% 47.2% 50.6% 43.4%
-IPI-M (Greshake et al., 2023) 11.1% 29.6% 4.3% 46.8% 20.8% 37.8% 33.2% 26.2%
+C-GCG [63] 43.5% 53.8% 55.4% 33.1% 20.4% 47.2% 50.6% 43.4%
+IPI-M [22] 11.1% 29.6% 4.3% 46.8% 20.8% 37.8% 33.2% 26.2%
 TPIA (ours) 85.8% 97.9% 78.5% 81.6% 70.5% 94.4% 97.4% 86.6%
 CodeGeeX2-6b No Attack 0.1% 13.4% 12.9% 19.0% 24.0% 21.9% 14.3% 15.0%
-C-GCG (Zou et al., 2023) 42.9% 46.1% 44.9% 41.7% 32.5% 51.7% 33.7% 41.9%
-IPI-M (Greshake et al., 2023) 9.5% 11.4% 5.3% 17.9% 0.9% 22.0% 11.5% 11.2%
+C-GCG [63] 42.9% 46.1% 44.9% 41.7% 32.5% 51.7% 33.7% 41.9%
+IPI-M [22] 9.5% 11.4% 5.3% 17.9% 0.9% 22.0% 11.5% 11.2%
 TPIA (ours) 79.6% 95.0% 80.8% 83.3% 72.0% 97.9% 85.6% 84.9%
 Malicious Objectives and Code Datasets.
-We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) (Top, 2024) across 3 programming languages.
+We hereby collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages: Python, Java, and C/C++.
 The straightforward objectives include using MODE_ECB for encryption (dubbed ‘ST0’), using ssl3 protocol (dubbed ‘ST1’), and executing “rm -rf /” (dubbed ‘ST2’).
-ST0 and ST1 are studied in classical code backdoor attacks (Schuster et al., 2021), while ST2 is a famous destructive instruction.
-CWE vulnerabilities are generally selected for experiments in code model research (He and Vechev, 2023; Yan et al., 2024), and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA).
-We consider three mainstream programming languages: Python, Java, and C/C++ to design the threat case for each objective.
-All the cases are constructed by the method described in Section 4.2 based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2).
-We will discuss the cases in detail in Appendix A.2.
-As shown in Table 1, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases.
-In addition, we collect corresponding code completion datasets, including Humaneval (Chen et al., 2021), Humaneval-x (Zheng et al., 2023), CodeXGLUE (Lu et al., 2021), MBPP (Austin et al., 2021), and Eval-Plus (Liu et al., 2024).
+ST0 and ST1 are studied in classical code backdoor attacks [46], while ST2 is a famous destructive instruction.
+CWE vulnerabilities are generally selected for experiments in code model research [23, 56], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA).
+We construct cases using the method described in Section IV-B based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2).
+As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases.
+In addition, we collect corresponding code completion datasets, including Humaneval [14], Humaneval-x [61], CodeXGLUE [38], MBPP [10], and Eval-Plus [35].
 The items in these datasets contain prompts and corresponding outputs across different code completion tasks, which can simulate unpredictable users’ instructions and context in real-world scenarios.
-Currently, most code completion datasets only cover Python, as shown in Table 1.
+Some datasets do not cover Java or C/C++, as shown in Table I.
 As a result, we have 7 sub-datasets (i.e., the subsection of a dataset containing a particular programming language) containing more than 14,000 data items totally.
 Model Selection.
-In our experiment, we select three models from two representative Code LLM series, including CodeGemma (Team et al., 2024) and Code-GeeX (Zheng et al., 2023), for evaluation.
-CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama (Roziere et al., 2023) on code completion tasks.
+In our experiment, we select three models from two representative Code LLM series, including CodeGemma [49] and Code-GeeX [61], for evaluation.
+CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [44] on code completion tasks.
 We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales.
-CodeGeeX series have corresponding applications (cod, 2024) with more than 535k downloads (on VScode (VSC, 2024) Apps shop).
+CodeGeeX series have corresponding applications [2] with more than 535k downloads (on VScode [6] Apps shop).
 We select CodeGeeX2-6b, a version with stable performance.
-We exclude earlier small-scale models, such as CodeBERT (Feng et al., 2020) and CodeT5 (Wang et al., 2021), because they perform much worse than large-scale models in code completion tasks.
+We exclude earlier small-scale code-oriented models, such as CodeBERT [18] and CodeT5 [51], because large-scale code-oriented models perform much better than them in code completion tasks.
 Baseline Selection.
 We compare our TPIA with GCG-based and manually designed IPI attacks.
-In particular, the vanilla GCG (Zou et al., 2023) is a representative PI attack and does not meet the requirements of our attack scenarios.
+In particular, the vanilla GCG [63] is a representative PI attack and does not meet the requirements of our attack scenarios.
 Therefore, we adapt GCG into an IPI attack that is applicable to Code LLMs, dubbed ‘C-GCG’ (code-oriented GCG).
-The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work (Greshake et al., 2023).
-IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments (as shown in Figure 11).
-Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have some inherent insecurity factors.
-In particular, we exclude backdoor attacks and adversarial attacks as our baselines because it is difficult to fairly compare these methods with TPIA.
-Specifically, backdoor attacks require control over the model training process, which is an overly strong assumption for our threat model.
-Adversarial attacks, on the other hand, cannot achieve target-specific attacks, leading to unreliable threat levels.
-We provide more explanations in the experiments in Appendix A.3.
+The manually designed IPI method (dubbed ‘IPI-M’) is designed based on the existing work [22].
+IPI-M limits the number of injected tokens to make the perturbation concise since the existing IPI attack usually uses a large number of readable comments.
+Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor.
 Evaluation Metrics.
-In this paper, we use attack success rate (ASR) to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness.
+In this paper, we use ASR to evaluate the effectiveness, not bad cases rate (NBR) and selective threat factor (STF) to evaluate the flexibility, and number of injected tokens and chars (NIT and NIC) to evaluate the stealthiness.
 For ASR, we consider an attack successful only when the generated code has equivalent functionality to the target code and appears in the correct position.
@@
 The more cases with sufficiently high ASRs (above a threshold $r_{2}$) and larger overall ASR value, the greater the upper limit of the attack’s threat.
-Specifically, NBR and STF can be formulated as follows:
-(7) NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$
+Their formulations are as follows:
+NBR $=1-\frac{\text{\# of cases where ASR}<r_{1}}{\text{Total \# of cases}},$ (8)
 STF $=\frac{\sum_{i\in\text{cases where ASR}>r_{2}}\text{ASR}_{i}}{\text{Total \# of cases}},$
-where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation. More details about the selection of metrics and values are in Appendix A.2.
+where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation.
 Hyperparameters and Computational Facilities.
@@
 During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100.
-For forward reasoning enhancement in Section 4.4, the number of tokens reinforced is set to $h=2$, as defined in Equation 6.
+For forward reasoning enhancement in Section IV-D, the number of tokens reinforced is set to $h=2$, as defined in Equation 6.
 For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10.
@@
 In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b.
-These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes.
-5.2. Evaluation on Open-source Code LLMs
+These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes across all the experiment threat cases.
+V-B Evaluation on Open-source Code LLMs
 We first perform the attack evaluation over open-source Code LLMs.
@@
 For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12.
-Since IPI-M needs sufficient tokens to be effective, we do not set a fixed length for this attack but limit the number of used tokens under 30 (details of IPI-M are described in Appendix A.2).
-Table 2 provides the average ASR for four methods.
+Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens.
+Table II provides the average ASR for four methods.
 The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
@@
 However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability.
-On the other hand, the worst-performing CodeGemma-2b is more easily affected by simpler and more direct baseline methods such as C-GCG and IPI-M; Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content.
+Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content.
 In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M.
@@
 In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness.
-According to our subsequent experiments in Section 5.3 and Section 5.4, increasing the length of TPIA or Top-$k$ value can achieve better attack performance.
-An interesting phenomenon is that IPI-M can sometimes enhance the security of the model’s output (producing a lower ASR than ‘No Attack’).
-We speculate that this is mostly because, in some cases, the generated perturbation contains information on the potential positions where vulnerabilities might occur but does not clearly convey the content of the malicious behavior to Code LLMs.
-As such, Code LLMs actually strengthen their reasoning about the locations of potential weaknesses and produce more secure outputs.
-It suggests that (parts of) our TPIA (e.g., those used in IPI-M) could potentially be used for enhancing the security of Code LLMs, by automatically optimizing the unfunctional parts of reference code to improve the security of the generated code.
-We will further explore it in our future work.
-Table 3. The attack evaluation results of NBR and STF against open-source Code LLMs.
+According to our subsequent experiments in Section V-C and Section V-D, increasing the length of TPIA or the Top-$k$ value can effectively enhance the attack performance of TPIA.
+TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs.
 Method Metric C/C++ Python Java
@@
 We use NBR to evaluate the attack flexibility for the first situation, and STF for the second situation.
-As shown in Table 3, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible.
+As shown in Table III, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible.
 Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat.
@@
 We hereby evaluate the trade-off between effectiveness and stealthiness.
-As we discussed in Section 3.3, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals.
+As we discussed in Section III-C, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals.
 Arguably, human inspection is currently the only feasible approach to detect TPIA attacks.
@@
 As shown in Figure 6, ASR and STF show a clear increase as NIT increases.
-When NIT is above 12, ASR and STF completely overlap, indicating that ASR exceeds 80% across all cases.
 Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values.
@@
 When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation).
-This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods, adding just a few tokens can be highly effective.
+This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods [22], adding just a few perturbation tokens can be highly effective.
 Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR.
@@
 (b) NIT/NIC under Different ASR
-Figure 6.
-The study of stealthiness.
+Figure 6: The study of stealthiness.
 We plot the above two graphs using NIT and ASR as the x-axis respectively.
-We sampled 11 different NITs, namely NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
-5.3. Evaluation on Commercial Code LLMs
+We sampled 11 different NITs, NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
+V-C Evaluation on Commercial Code LLMs
 In previous experiments, we have demonstrated that TPIA reaches sufficient effectiveness, flexibility, and stealthiness on open-source models.
@@
 A more challenging situation is when the attacker knows nothing about the model used by the target application, or the model itself is closed-source (such as CodeX).
-This is precisely the case with GitHub Copilot.
-We simulate this scenario by using perturbations generated by CodeGeeX2-6b to attack GitHub Copilot.
-Table 4. The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
+We simulate this scenario by attacking GitHub Copilot (supported by CodeX).
+To align with gray-box setting, we adopt CodeGeeX2-6b to generate malicious perturbations.
+TABLE IV: The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
 Scenario Language ASR NBR STF
@@
 (b) Target Completion
-Figure 7.
-The example of a successful attack in a real-world black-box scenario (Github Copilot) across files.
-The perturbation is inserted into the comment of a related file and influence the suggestion of LLM in main file.
+Figure 7: The example of a successful attack in a real-world black-box scenario (Github Copilot) across files.
+The perturbation is inserted into the comment of a function.
+When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after TRG).
 Settings.
 We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total).
-Other settings are the same as those described in Section 5.2.
+Other settings are the same as those described in Section V-B.
 Results.
-As shown in Table 4, in the gray-box scenario, TPIA still demonstrates great attack performance.
+As shown in Table IV, in the gray-box scenario, TPIA still demonstrates great attack performance.
 Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively.
@@
 In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility.
-Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to real-world commercial applications.
+Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
 In Figure 7, we present a representative case of ST2 under the black-box scenario.
 The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function.
-The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction.
-5.4. Ablation Study
+The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
+V-D Ablation Study
 In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for TPIA attacks.
@@
 (b) STF Study
-Figure 8.
-Ablation study of keywords number and Top-$k$.
+Figure 8: Ablation study of keywords number and Top-$k$.
 We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords.
@@
 (b) FRE
-Figure 9. Ablation study of FRE. We compare the differences in ASR and loss function values both with and without using FRE. 1-key1 and 1-key2 represent keeping one of the two keywords, respectively.
+Figure 9: Ablation study of FRE.
+We compare the differences in ASR and loss function values both with and without using FRE module.
+1-key1 and 1-key2 represent keeping one of the two selected keywords, respectively.
 Effects of Keyword-based Designing.
@@
 Firstly, under the same conditions, using FRE can lead to a higher ASR.
-Secondly, without FRE, the attack success rate is still very low ($<20\%$) even when the loss value is lower than that with FRE.
-Thirdly, FRE has a greater impact on attack success rate than keyword-based design.
+Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE.
+Thirdly, FRE has a greater impact on ASR than keyword-based design.
 These results verify the necessity of forward reasoning enhancement in our TPIA method.
-6. Discussions
-6.1. Resistance to Potential Defenses
-In Section 3.3, we have explained why our attack can fundamentally circumvent existing malicious code detection methods.
+VI Discussions
+VI-A Resistance to Potential Defenses
 In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack.
-In general, we can roughly categorize these methods into detection and removal and design following potential defense methods.
+In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against TPIA.
+Figure 10: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b.
+The two types are almost indistinguishable.
 Detection: Semantic Analysis.
 Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code.
-Since detection methods based on code functionality are ineffective against non-functional perturbations, it is necessary to adopt a semantic analysis approach.
-This involves examining the differences between the benign code and the suspicious code in the embedding layer of Code LLMs.
+Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach.
+This involves examining the differences between the benign code and the injected code in the embedding layer of Code LLMs.
 However, this method may be difficult to implement.
 We find that code segments injected with perturbations might not have obvious semantic differences from benign ones.
-Specifically, as shown in Figure 16, we construct classification prompts and extract their embedding layer features on Code LLMs, finding that the two types of code segments are almost completely indistinguishable.
-We present more details and analyses in Appendix A.3.
-Figure 10.
-The resistance to the removal of chars.
+Specifically, as shown in Figure 10, we sample 1640 code segments and construct classification prompts.
+The results show that benign code and injected code are almost completely indistinguishable.
+Figure 11: The resistance to the removal of chars.
 We remove characters from the perturbation, selecting 5% for removal each time, and we calculate the average ASR for each programming language across applicable cases.
@@
 This type of defense involves removing perturbations by altering the non-functional parts of the code after introducing external code.
-In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect malicious code segments (as mentioned above).
-For example, if preserving code readability is a priority, we cannot remove all perturbations, allowing only for a lower degree of modifications.
+In general, it entails a trade-off between code readability and the strength of removal since we cannot faithfully detect injected code segments (as mentioned above).
 To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations.
-As shown in Figure 10, if the removal rate of perturbations cannot be ensured to be above 0.5, TPIA still maintains a sufficiently high ASR.
-On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table 5.
+As shown in Figure 11, TPIA still maintains a sufficiently high ASR under 0.5 removal rate.
+On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table V.
 Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
-Table 5. ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments form.
+TABLE V: ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments.
 […]
@@
 Java 66.5% -9.0%
-6.2. Potential Limitations and Future Directions
-Although TPIA is effective in attacking the evaluated LLMs and commercial applications, we have to admit that it still has some drawbacks.
+VI-B Potential Limitations and Future Directions
+TPIA still has some drawbacks.
 Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords.
@@
 Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent.
-We believe that readable and benign-looking perturbations may pose a more realistic threat and represent an important direction for future research.
-7. Conclusion
+In future work, we will explore how to use LLM as an optimizer to further improve this.
+VII Conclusion
 In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks.
@@
 We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
-Our work alerted a new serious security threat in using Code LLMs in practice.
-Appendix A Appendix
-A.1. Differences with Existing Methods
-GCG attack.
-In the domain of generative models, there exists some adversarial prompt injection attack methods, such as universal trigger (Wallace et al., 2019) and well-known GCG (Zou et al., 2023).
-Taking GCG as an example, the adversarial goal of the attacker is to force LLMs to generate affirmative responses beginning with “Sure, here’s…,” thereby disrupting security alignment (Ji et al., 2024).
-Unlike GCG, TPIA emphasizes the feasible usage of perturbations and performance during indirect attacks.
-First, TPIA transfers different malicious objectives to perturbations to induce specific code generation, while GCG aims to generate the most powerful perturbation to break the safety guard.
-Second, TPIA considers a scenario where the users are victims and uses keyword-based perturbation to save the number of perturbation tokens, while GCG assumes that the users of LLMs are malicious attackers and allow unlimited tokens.
-Third, TPIA proposes a code-oriented task construction framework and forward reasoning enhancement technique to ensure effectiveness and robustness in uncertain contexts, while GCG has a controllable context and fixes the perturbation at the end of the prompt.
-What is more, straightforward application of adversarial perturbation methods like GCG on code completion attack has a terrible performance.
-As shown in Figure 9, 10-adv attacks with “0-keyword no FRE” refers to the 0% ASR.
-Codebreaker and SVEN.
-In the domain of code, the newly proposed backdoor attack method, Codebreaker (Yan et al., 2024), also leverages the feedback of LLMs for optimization.
-However, unlike TPIA, Codebreaker emphasizes using the coding capabilities of LLMs to reduce the probability of the malicious payload code being detected, thereby increasing the success rate of backdoor attacks, without removing the reliance on the poisoning process.
-A method called SVEN (He and Vechev, 2023) involves purposefully generating high-risk code.
-However, this method requires Prefix-tuning on Code LLMs, adding new parameters to the original model parameters, making it unsuitable as an attack method.
-Additionally, when generating high-risk code, SVEN cannot control the content of risk code, nor can it achieve target-specific attacks.
-A.2. Additional Details
-Figure 11.
-Greshake et al.
-(Greshake et al., 2023) use a large amount of disruptive information and explicit malicious code snippets in the comments, which has a significant impact on the feasibility of the attack.
-Baseline Details.
-IPI-M and C-GCG are important baselines in our effectiveness experiment.
-IPI-M uses injection instructions in natural language form to interfere with the model’s reasoning results, commanding the model to generate insecure code.
-The original IPI-M places a large amount of disruptive commentary along with the malicious code to be inserted in the comment section, as shown in Figure 11.
-We argue that this method has very low feasibility.
-Despite the attack path of IPI does not need to worry about being detected by malicious code detection methods, there is a risk of being discovered and removed once the victim becomes aware and actively pays attention to IPI attack, especially perturbations with poor visual stealthiness.
-Therefore, we attempt to simplify IPI-M into concise and clear template-based injection sentences, as shown in the table.
-C-GCG mainly follows the GCG method, using an adversarial prefix to achieve malicious purposes.
-At the same time, C-GCG does not adopt the two optimization methods from Section 4.4 and Section 4.5.
-C-GCG performs much worse than TPIA in the main experiments.
-Value and Metric Selection.
-We select the values of $r_{1}$ and $r_{2}$ based on the attack’s performance in the experiment.
-Specifically, in our experiments, 20% ASR is smaller than the ASR of no attack baseline in some situations, and 80% ASR outperforms the best performance of all the baselines.
-In addition, we find that only a few subcases (determined by the model, dataset, attack method, and case) have a success rate between 20% and 80%, accounting for 7.5%.
-Therefore, setting 20% and 80% as threshold of bad case and outstanding case respectively is appropriate.
-As for metrics, we have already explained why metrics such as NBR and STF are used in the main experiments.
-Here we further explain some details related to them.
-First, NBR does not take the specific ASR value into account, as this metric is used to measure the diversity of attacks, rather than ensuring stable success rates.
-Second, STF targets the situation where attackers aim to cause as much damage as possible, and it varies with differences in ASR.
-In other words, even a difference between 100% ASR and 95% ASR can have a real impact on the overall threat level.
-Therefore, the calculation of STF considers the specific ASR value.
-In addition, we do not use common metrics for code generation tasks, such as BLEU or CodeBLEU (Ren et al., 2020), because they do not directly reflect the effectiveness of the attacks.
-Overall, we attempt to confirm the realistic meaning of each evaluation metric from different perspectives.
-Case Study.
-Here we show the details of our malicious objectives and threat cases.
-As shown in Table 6, we design 13 malicious objectives and 31 threat cases in 3 programming languages based on CWE top-10 KEV, exisiting research and famous destructive instruction.
-In programming process, minor changes in the code can cause significant functional alterations.
-We roughly divide the causes of vulnerabilities in all malicious objectives into three categories: skipping critical process (CWE20, 78, 502, 918, 22, 306), wrong memory usage (CWE416, 122, 787, 843), and outputting dangerous snippets (ST0, 1, 2).
-All these vulnerabilities result in severe consequences.
-Table 6. Description of Threat Cases.
-Case Description From
-CWE416 Use after free. CWE 1st KEV
-CWE122 Heap-based buffer overflow. CWE 2nd KEV
-CWE787 Out-of-bounds write. CWE 3rd KEV
-CWE20 Improper Input Validation CWE 4th KEV
-CWE78 ‘OS Command Injection’ CWE 5th KEV
-CWE502 Deserialization of untrusted data. CWE 6th KEV
-CWE918 Server-side request forgery. CWE 7th KEV
-CWE843 Access of resource using incompatible type. CWE 8th KEV
-CWE22 ‘Path Traversal’ CWE 9th KEV
-CWE306 Missing authentication for critical Function. CWE 10th KEV
-ST0 Use ECB mode to encrpty. Schuster et al.(Schuster et al., 2021)
-ST1 Use SSL3 protocol Schuster et al.(Schuster et al., 2021)
-ST2 Executing “rm -rf / Famous destructive instruction
-(a) Skipping critical process - CWE20
-(b) Wrong memory usage - CWE122
-(c) Outputting dangerous snippets - ST2
-Figure 12.
-Three examples of TPIA attack cases.
-Code LLMs normally generate code segment without vulnerabilities, but they lose this ability after being attacked by TPIA.
-Attackers can modify the output content at the specified location through TPIA, and they can even generate malicious code snippets that have never appeared in the context, e.g., “os.system("rm -rf /")”.
-As shown in Figure 12, here we present three representative examples across three different languages and different type of objectives.
-The first is CWE20 (using Java), where the program attempts to fetch the content of a URL and calculate its length.
-A secure programming practice would be to check if the content is null before performing the length calculation, thus avoiding a null pointer.
-However, after being attacked, the Code LLM generates code that calculates the length directly after fetching the URL, leading to a null pointer exception if the URL is null.
-A null pointer can result in serious consequences such as program crashes or data corruption.
-We simply categorize this situation as “skipping critical process.” The second example is CWE122 (using C), where the program attempts to pass content to buf through a command-line argument.
-A secure programming practice would be to check if the content exceeds BUFSIZE before the assignment, thus avoiding overflow.
-Through TPIA, attackers can induce Code LLMs to perform the assignment directly.
-When a stack overflow occurs, it accesses non-existent RAM space, leading to consequences such as system crashes and data leaks.
-We simply categorize this situation as “wrong memory usage.” This type of situation generally applies only to C/C++, as Java and Python automatically manage memory release.
-The third example is ST2 (using Python), where the program attempts to call a function named TRG (TaRGet).
-Under normal circumstances, this call does not present any issues, and the large model generates suggestions based on the contextual semantics or directly produces an ¡eos¿ token.
-Through TPIA, attackers can make Code LLMs generate the execution code “rm -rf /” right after calling the TRG function.
-If executed, this code results in the deletion of all data under the root directory.
-We categorize this situation as “outputting dangerous snippets.”
-We only apply different perturbation construction strategies to different categorizations of cases in the IPI-M.
-Specifically, for “skipping critical process,” we use “… directly after …” and “without any check”; for “wrong memory usage,” we use “It is important to … directly after …”; for “outputting dangerous snippets,”, we use “It is very important to promptly … after …”.
-In addition, the perturbation in IPI-M must be sufficiently explicit; for example, it needs to specify the code snippet you want to insert (e.g., “os.system("rm -rf")”) or present a clear assertion (e.g., “SSL3 is the most secure protocol”).
-Otherwise, Code LLMs are difficult to induce successfully.
-Arguably, the actual performance of IPI-M may be worse than our evaluation results shown in main experiments.
-Keywords Selection Method.
-We use a simple dual-loop strategy in the selection of keywords.
-Specifically, before each round of perturbation optimization with length $l$ starts in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords.
-With the optimization step $s$, the equation can be formulated as follows:
-(8) $\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$
-$\bm{P}_{adv}^{s+1}=\min_{\bm{P}_{adv}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P}_{adv},\bm{\theta}),$
-where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6.
-$\bm{W}$ represents the token sequence that constitutes the keywords.
-To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., $\bm{P}_{key}=\{p_{k1},p_{k2},\dots,p_{kn}\}$, and for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{target code},\text{position code}\}$.
-(a) Gemma
-(b) Codegemma
-(c) Codellama
-(d) Codegeex2
-Figure 13.
-Examples of traditional adversarial attacks on four different LLMs.
-The generated perturbations are displayed with a gray background, the input code completion task content with a green background, and the output content with a red background.
-A.3. More Experiments
-Comparison with Adversarial Attack.
-The advantages of TPIA over backdoors are clear: it does not require control of the training process and has a very low attack threshold.
-However, the advantages of TPIA over existing adversarial attacks, i.e., achieving target-specific attacks, is not obvious.
-Figure 14.
-The ASR across different cases.
-The experimental results shown in this figure are the average values calculated from the effectiveness experiments involving 3 programming languages, 4 datasets, and methods excluding Clean Accuracy.
-Here, we propose an imperfect but highly intuitive scoring method to compare TPIA with state-of-the-art adversarial attack method (Jha and Reddy, 2023).
-We divide the output results into three levels: (1) generating code with normal functionality, with a harmfulness score of 0 (level 0); (2) generating code that cannot compile or has functional errors, with a harmfulness score of 0.5 (level 1); (3) generating code with stealthy risk, with a harmfulness score of 1 (level 2).
-In terms of the resulting effects, the second level will lead to a decrease in the usability of Code LLM, and the third level not only affects the usability but also causes unpredictable financial losses to the users of the Code LLMs.
-We evaluate the performance on three cases: S0, S1, and S2, using four models: Gemma-7b, CodeGemma-7b, CodeLlama-7b, and CodeGeeX-6b (Due to the significant differences in resistance to traditional adversarial attacks across different models, we use more types of LLMs than in the main experiments, including a general-purpose LLM).
-Figure 15.
-A comparison of the harmfulness between CodeAttack (CA) and TPIA.
-The final scores were normalized to the 0-100.
-We use light colors (CA lv1 and TPIA lv1) to represent ratings from level 1 harmfulness and dark colors (CA lv2 and TPIA lv2) to represent ratings from level 2 harmfulness.
-Figure 16.
-The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b.
-The two types are almost indistinguishable.
-As shown in Figure 15, TPIA significantly outperforms CodeAttack in scores.
-This is because TPIA consistently achieves level 2 harmfulness, whereas CodeAttack tends to reach level 1.
-In fact, the final result of CodeAttack is not invalid: the quality of output is greatly impacted, as shown in Figure 13.
-However, the output code usually cannot be correctly compiled.
-In other word, this type of attack is unlikely to threaten the user’s experience, although it successfully reduces the usability of Code LLM and has a negative impact on its commercial applications.
-In contrast, TPIA also impairs the usability of the victim Code LLMs, and can cause terrible subsequences for victim users.
-In summary, TPIA is capable of conducting more harmful attacks with a higher success rate, making it more suitable for LLM-based code completion scenarios compared to traditional adversarial attack methods.
-Effectiveness across Different Cases.
-In the main experiments, we have validated the outstanding attack performance of TPIA from three perspectives: model, dataset, and programming language.
-Here, we supplement the attack details for different attacks on a single case.
-Specifically, we use ASR to evaluate the performance of TPIA and the baselines across different cases.
-We present the specific average ASR for each case.
-As shown in Figure 14, TPIA achieves a certain level of effectiveness across all cases, with an overall average ASR of 83.4%.
-However, C-GCG and IPI-M both have cases where they are completely ineffective, especially for IPI-M.
-This demonstrates that directly using traditional GCG methods or manually designing perturbations is not effective in some scenarios.
-We believe this involves three reasons: (1) The baseline methods are not fully suitable for contexts with structured features, which is a fundamental characteristic of the code completion task; (2) The specific malicious objective include complex semantics such as target code, position code, and “disregarding security attributes,” which are sometimes not fully captured by the perturbations generated by the baseline methods; (3) The impact of keyword-based designing and FRE; our ablation study proves that these two modules play a critical role in TPIA’s effectiveness.
-According to the results, the most challenging to achieve is the straightforward case ST2.
-This case requires generating an obviously malicious command in a completely unrelated context.
-Another difficult case to implement is CWE22 “Path Traversal,” which involves using unconventional paths like “../” to access files that are normally inaccessible.
-This case is relatively easier to achieve in Python, where the Code LLMs tend to suggest directly reading the path.
-However, it is more challenging to implement in Java and C/C++, as Code LLMs often only recommend checking the path rather than directly reading the file.
-CWE843 involves manipulating union members in a struct in C/C++.
-This represents a more complex malicious objective, where we attempt to manipulate one value within the union using an untrusted value.
-This is an unconventional operation that is difficult to describe using natural language, making it challenging for IPI-M to successfully attack.
-In contrast, methods like TPIA and C-GCG, which convey malicious instructions through token selection, are able to achieve success.
-Arguably, TPIA has a clear advantage in attack performance for all the cases we have collected.
-Semantic Analysis on Payload Code.
-In this experiment, we attempt to detect malicious payload code by semantic analysis.
-We observe the classification features of code segment in the embedding layer of CodeGemma-2b.
-Specifically, we add the classification prompt keyword [CLS] to 820 original code snippets and 820 malicious payload code snippets totally.
-and then input them into the model to analysis their semantic features.
-We examine the t-SNE (Van der Maaten and Hinton, 2008) distribution of the classification keyword [CLS] in the embedding layer, which is a widely used method for classification tasks (Abdelnabi et al., 2024).
-As shown in Figure 16, the features of the two types of text are almost completely intertwined, making it nearly impossible to distinguish between the injected code data and the original code data.
-In addition to observing the characteristics of layer 17 shown in Figure 16, we also examine the data features of layers 1 through 16.
-There is no significant difference observed in these layers compared to layer 17.
-There are several reasons for this situation.
-First, code data, including comments, tend to have lower fluency and coherence with high token complexity, which poses a significant challenge for semantic analysis tasks.
-Second, the perturbations used by TPIA occupy an extremely small number of tokens, exerting only a minor influence on the overall semantics of the injected data, making them hard to detect.
-Third, the semantic impact of TPIA differs from traditional injection attacks in that it does not require an extremely high intensity of information to counteract the model’s secure alignment; instead, it guides the model to generate vulnerabilities that could occur naturally, thus making detection through semantic analysis difficult.
-Our experiment results indicate that simple semantic analysis methods are not effective in defending against TPIA.
-This indicates that detecting code snippets containing malicious perturbations generated by TPIA through semantic analysis is not an effective approach.
+Our work alerted a new serious security threat in using Code LLMs.
+Acknowledgment
+We would sincerely thank Prof. Kui Ren and Prof. Chun Chen from the Zhejiang University for their helpful comments and suggestions on an early draft of this paper.

### v6

@@
-TPIA: Towards Target-specific Prompt Injection Attack against Code-oriented LLMs
+ShadowCode: Towards (Automatic) External Prompt Injection Attack against Code LLMs
 Abstract
-Recently, code-oriented large language models (Code LLMs) have been widely and successfully exploited to simplify and facilitate programming.
-Unfortunately, a few pioneering works revealed that these Code LLMs are vulnerable to backdoor and adversarial attacks.
-The former poisons the training data or model parameters, hijacking the LLMs to generate malicious code snippets when encountering the trigger.
-The latter crafts malicious adversarial input codes to reduce the quality of the generated codes.
-In this paper, we reveal that both attacks have some inherent limitations: backdoor attacks rely on the adversary’s capability of controlling the model training process, which may not be practical; adversarial attacks struggle with fulfilling specific malicious purposes.
-To alleviate these problems, this paper presents a novel attack paradigm against Code LLMs, namely target-specific prompt injection attack (TPIA).
-TPIA generates non-functional perturbations containing the information of malicious instructions and inserts them into the victim’s code context by spreading them into potentially used dependencies (e.g., packages or RAG’s knowledge base).
-It induces the Code LLMs to generate attacker-specified malicious code snippets at the target location.
-In general, we compress the attacker-specified malicious objective into the perturbation by adversarial optimization based on greedy token search.
-We collect 13 representative malicious objectives to design 31 threat cases for three popular programming languages.
-We show that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
-I Introduction
-Recently, generative artificial intelligence (GenAI) techniques have been widely and successfully exploited in code engineering, such as automated programming and code generation.
-In particular, code-oriented large language models (Code LLMs), e.g., CodeX [53], CodeGemma [49], and CodeGeex [61], have demonstrated immense potential in code completion [^1] Although there are other programming tasks such as code translation and code summarization, code completion remains the mainstream research focus [60, 34] and the fundamental paradigm in commercial practice [19, 2]..
-They can efficiently generate complete functional code based on incomplete code snippets and text prompts.
-Driven by such attractive capability, numerous applications integrated with Code LLMs have been released, such as Github Copilot [19] developed jointly by Github [3] and OpenAI [4].
-These applications can provide real-time code completion suggestions in the Integrated Development Environment (IDE), significantly improving programming efficiency.
-Despite their remarkable performance, recent studies revealed that Code LLMs are vulnerable to some well-designed attacks, which maliciously manipulate the functional execution of Code LLMs in practice.
-In general, existing attacks can be divided into two main categories.
-(1) Backdoor attacks [46, 29, 57].
-They intend to compromise the models by poisoning the training data [46, 57] or model parameters [29], such that the infected models will generate targeted code snippets, when the input contains attacker-specified trigger patterns.
-They can achieve diverse attack consequences, such as incorrect code additions, deletions or modifications [29], or implementation of less secure encryption rules or protocols [46].
-(2) Adversarial attacks [62, 12, 24].
-They aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs.
-For example, Jha et al.[24] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs.
-However, we reveal that existing attacks suffer from some inherent limitations, hindering their realistic threats to some extent.
-Specifically, backdoor attacks can achieve specific malicious objectives but require the capability of interfering the model training, e.g., modifying some training samples or even model parameters.
-In real-world scenarios, developers of Code LLMs or integrated applications usually obtain the models or datasets from trusted entities (e.g., renowned research institutes or companies).
-This mitigates the backdoor threats from the source.
-Besides, decoupling the implantation of the malicious payload and backdoor trigger from poisoning operations is highly difficult.
-Arguably, without controlling the training process of Code LLMs, backdoors cannot exist; On the other hand, existing adversarial attacks [12, 62, 24] can hardly achieve precise malicious objectives, although they do not require intervention in the training process.
-This is mostly due to their particular goal of attackers, i.e., finding adversarial samples that can degrade the quality of code generation, can inevitably weaken Code LLMs’ understanding of the input.
-Driven by these findings, an intriguing and critical question arises: Has the application of Code LLMs already achieved a sufficient level of security in practice?
+Recent advancements have led to the widespread adoption of code-oriented large language models (Code LLMs) for programming tasks.
+Despite their success in deployment, their security research is left far behind.
+This paper introduces a new attack paradigm: (automatic) external prompt injection against Code LLMs, where attackers generate concise, non-functional induced perturbations and inject them within a victim’s code context.
+These induced perturbations can be disseminated through commonly used dependencies (e.g., packages or RAG’s knowledge base), manipulating Code LLMs to achieve malicious objectives during the code completion process.
+Compared to existing attacks, this method is more realistic and threatening: it does not necessitate control over the model’s training process, unlike backdoor attacks, and can achieve specific malicious objectives that are challenging for adversarial attacks.
+Furthermore, we propose ShadowCode, a simple yet effective method that automatically generates induced perturbations based on code simulation to achieve effective and stealthy external prompt injection.
+ShadowCode designs its perturbation optimization objectives by simulating realistic code contexts and employs a greedy optimization approach with two enhancement modules: forward reasoning enhancement and keyword-based perturbation design.
+We evaluate our method across 13 distinct malicious objectives, generating 31 threat cases spanning three popular programming languages.
+Our results demonstrate that ShadowCode successfully attacks three representative open-source Code LLMs (achieving up to a 97.9% attack success rate) and two mainstream commercial Code LLM-integrated applications (with over 90% attack success rate) across all threat cases, using only a 12-token non-functional induced perturbation.
+The code is available at https://github.com/LianPing-cyber/ShadowCodeEPI.
+Index Terms:
+Code-oriented Large Language Models, Prompt Injection Attack, Code Security, AI Security
+1 Introduction
+Recent advancements in generative artificial intelligence (GenAI) have led to its widespread and successful application in code engineering, such as automated programming and code generation [1, 2].
+Notably, code-oriented large language models (Code LLMs) such as Codex [3], CodeGemma [4], and CodeGeex [5] have shown significant promise in code completion [^1] Code completion currently remains a primary research focus [6, 7] and a foundational paradigm in commercial practice [1, 2], despite the existence of other programming tasks (e.g., code summarization).
+These models efficiently generate complete, functional code from incomplete snippets and text prompts.
+This compelling capability has spurred the development of numerous Code LLM-integrated applications, such as Github Copilot [1] developed jointly by Github [8] and OpenAI [9].
+Such applications enhance programming efficiency by offering code completion suggestions directly within Integrated Development Environments (IDEs).
+Despite their impressive performance, recent studies have shown that Code LLMs are susceptible to sophisticated attacks that maliciously manipulate their functional execution.
+These attacks generally fall into two primary categories: (1) Backdoor attacks [10, 11, 12], aim to compromise models by poisoning training data [10, 12] or modifying model parameters [11].
+This manipulation causes infected models to generate targeted code snippets when specific, attacker-defined trigger patterns are present in the input.
+The consequences of such attacks can be diverse, including incorrect code additions, deletions, or modifications [11], as well as the implementation of less secure encryption rules or protocols [10].
+(2) Adversarial attacks [13, 14] focus on degrading the quality of the generated code by crafting adversarial examples.
+These examples are typically based on approximated gradients with respect to the model’s inputs and outputs.
+For instance, Jha et al.[14] demonstrated performance degradation in Code LLMs by altering the function and variable names.
+However, we contend that existing attacks possess inherent limitations, hindering their practical threat.
+Specifically, while backdoor attacks can achieve precise malicious objectives, they necessitate intervention in the model training process, such as modifying training samples [12, 15] or even altering model parameters [11].
+In real-world scenarios, developers of Code LLM or integrated applications usually acquire the models or datasets from trusted entities (e.g., renowned research institutes or companies), thereby mitigating backdoor threats from their origin.
+Conversely, existing adversarial attacks [13, 14], despite not requiring training intervention, struggle to achieve precise malicious objectives.
+This is primarily attributable to their core aim: identifying adversarial samples that degrade code generation quality, which inherently compromises the Code LLM’s understanding of the input.
+These findings raise an intriguing and critical question: Has the application of Code LLMs already achieved a sufficient level of security in practice?
 Unfortunately, the answer to the above question is negative.
-In this paper, we introduce a new attack paradigm against Code LLMs, dubbed target-specific prompt injection attack (TPIA) [^2] Note that our new attack paradigm (i.e., TPIA) significantly differs from existing prompt injection methods, e.g., universal trigger [50] and Greedy Coordinate Gradient (GCG) [63].
-Please find more details in Section II..
-TPIA leads to a more realistic threat by inheriting the advantageous features of both backdoor and adversarial attacks.
-In general, as shown in Figure 1, the TPIA attacker generates a concise non-functional perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion) , and inserts it into the victim’s code context.
-When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
-Arguably, attackers can easily introduce particular malicious non-functional perturbations to the victim’s code context in practice.
-For example, users may introduce external code by copying or installing packages; Attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27].
-In particular, different from supply chain attacks [41] that can be easily detected by static analysis [33] or sandboxing [9] since they directly insert malicious functional code to implant malicious behaviors, our TPIA can circumvent existing mainstream detection methods by introducing only non-functional code.
-Figure 1: The attack paragdim of our TPIA.
-A concise non-functional attacker-generated perturbation is injected into one of the project’s files.
-When the victim user completes the code via the Code LLM, the model will be induced by the incorporated attacker-specified code to generate an unsafe output at specified position, e.g., in edit_part() function.
-In general, our attack consists of two main stages.
-In the first stage, we construct the input and target output based on the attacker-specified objective.
-We represent the malicious objective as an output tuple (target position, target code), which aims to induce the Code LLM to generate the target code at the target position.
-We further expand the tuple into a complete task code, including a conditional code (e.g., necessary package references), context code (e.g., noises), and position code to simulate the actual attack scenario.
-In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective.
-Specifically, we obtain the base Code LLM and use adversarial gradient computation to search for the optimal perturbation tokens according to the likelihood of the Code LLM in generating the target code.
-This computation involves two important optimizations: (1) Forward reasoning enhancement: it allows the optimization direction of adversarial computation to be closer to the actual reasoning process of the Code LLM (forward token sampling with randomness), thus improving the attack’s effectiveness and robustness.
-(2) Keyword-based perturbation design: it offloads insensitive parts of the malicious objective into static tokens, significantly reducing the number of tokens required for a successful attack.
-We have collected 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [1] across 3 programming languages.
-There are a total of 31 threat cases since some of these malicious objectives apply only on C/C++.
-We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b, and CodeGeex2-6b.
-The results demonstrate that TPIA can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness.
-In addition, we evaluate TPIA on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 535k installations in the VScode [6] extension shop) and Github Copilot [19] (developed by Github and OpenAI [4]).
+Drawing inspiration from prompt injection (PI) attacks (e.g., jailbreak attacks) [16, 17], which leverage carefully designed prompts to force LLMs into generating malicious output, we observe that prompt perturbations introduce a further risk of malicious code completion.
+In this paper, we introduce a new attack paradigm: external prompt injection against Code LLMs [^2] We note that there are currently a few pioneering methods related to ours (generally dubbed ‘indirect prompt injection’) targeting Code LLMs [18, 19].
+However, these works only provide simple examples, and their injection content is manually designed.
+More details and discussions are described in Section 2.3..
+As illustrated in Figure 1, an attacker crafts a concise non-functional induced perturbation (e.g., comments) containing malicious instructions (e.g., generating code snippets for data deletion [20]).
+This perturbation is then inserted into the victim’s code context.
+When the victim requests the Code LLM to perform code completion based on this infected code context, the implanted perturbation will induce the model to respond to the particular malicious instructions at specific locations.
+Arguably, attackers can now easily and stealthily introduce particular malicious perturbations to the victim’s code context in practice.
+For example, users may introduce external code by copying or installing packages; attackers can also upload their malicious code to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded and used in a knowledge base to enhance the performance of Code LLMs [21] through retrieval-augmented generation [22].
+In particular, different from supply chain attacks [23], which are often detectable by static analysis [24] or sandboxing [25], external prompt injection introduces only non-functional code rather than malicious functional code to implant malicious behaviors, thereby circumventing mainstream detection methods.
+The attack scenarios of traditional PI differ from our external prompt injection.
+Specifically, conventional PI attacks are typically executed by malicious users directly targeting large language models (LLMs), whereas our focus is on a malicious third-party attacker aiming to compromise the users of a victim application.
+More importantly, simply using existing PI attacks to solve the problem in our setting is infeasible, as traditional PI attacks [26, 17] rely on strong assumptions.
+Specifically: (1) They assume controlled positioning, requiring the malicious instruction/perturbation to be placed at a fixed position within the model’s input (e.g., via prefix or suffix).
+(2) They usually assume an unlimited token budget, allowing for the use of hundreds of tokens and clear malicious instructions (e.g., “make a bomb”).
+These assumptions lead to insufficient attack effectiveness and stealthiness in practice.
+Figure 1: The attack paradigm of external prompt injection involves injecting a concise, non-functional, attacker-generated perturbation into a victim’s code context.
+When the victim user completes the code via a Code LLM, the model is induced by this incorporated attacker-specified code to generate an unsafe output at a specified position, e.g., in edit_part() function in this example.
+To address these challenges, we propose a greedy perturbation optimization method, named ShadowCode, for external prompt injection based on code simulation.
+ShadowCode generally comprises two main stages.
+In the first stage, we design the perturbation optimization objective based on the attacker-specified malicious objective.
+Specifically, we represent the malicious objective as an output tuple (position code, target code), which induces the Code LLM to generate the malicious target code immediately after the position code.
+Subsequently, we conduct code contextual simulation, which includes conditional code (e.g., necessary packages and method declarations), noise code (i.e., target-agnostic random code), as well as pre-defined position code and target code.
+This simulation aims to mimic diverse, real-world code snippets, thereby ensuring ShadowCode’s effectiveness in victim-specified, uncontrolled contextual environments.
+Finally, we design the loss function to identify the optimal induced perturbation that achieves this simulated outcome.
+In the second stage, we employ a greedy search to optimize the perturbation, incorporating two key designs: (1) forward reasoning enhancement: We increase the weight of the first few tokens in the target code in calculating the loss to enhance the effectiveness of ShadowCode under Code LLMs’ autoregressive generation manner; (2) keyword-based perturbation design: we introduce a dynamic prefix containing a few key tokens from the output tuple as the augmentation of induced perturbation during optimization, aiming to reduce the length of final induced perturbation and avoid optimization failure caused by greedy search.
+This design is motivated by the understanding that key tokens convey essential information from the malicious objective, thereby facilitating the desired induction.
+We collect 13 representative malicious objectives, including 3 straightforward yet representative malicious objectives and the Top-10 CWE Known Exploited Vulnerabilities (KEV) [27].
+These objectives are converted into specific cases across 3 programming languages, leading to a total of 31 threat cases since some of these malicious objectives apply only on C/C++.
+We evaluate our attack against various representative open-source Code LLMs, including CodeGemma-2b, CodeGemma-7b [4], and CodeGeex2-6b [5].
+The results demonstrate that ShadowCode can reach an 84.5% average attack success rate (ASR) over all cases and obtain over 80% ASR using only 8 tokens, indicating its effectiveness and stealthiness.
+In addition, we evaluate ShadowCode on mainstream commercial Code LLM-integrated applications, including CodeGeeX [2] (with more than 935k installations in the VScode [28] extension marketplace) and Github Copilot [1] (developed by Github and OpenAI [9]).
 These applications correspond to the gray-box and black-box attack scenarios, respectively.
-Our TPIA achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its practical threat in the real world.
-In summary, our main contributions are three-fold.
-  • 1)
-    We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs.
-  • 2)
-    We introduce a new attack paradigm (i.e., TPIA) against Code LLMs.
-It can induce the Code LLM to generate an attacker-specified malicious code snippet at specific locations by simply inserting a concise perturbation into the victim’s code context.
-  • 3)
-    We conduct extensive experiments to evaluate TPIA under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
-II Background and Related Work
-II-A Code LLMs for Code Completion
-Code and its comments can be considered as a corpus for deep learning models, enabling them to handle code tasks, such as code translation [48, 45], and code summarization [40, 8].
-The emergence of LLMs provides new alternatives to the coding tasks.
-By leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [53], CodeLlama [44], CodeGemma [49] and CodeGeex [61].
-These Code LLMs can provide detailed explanations about the intrinsic logic of code and easily generate a vast of functionally complex code [54].
-In this paper, we focus on code completion [34, 16, 60].
-This is a paradigmatic task that involves generating new code from incomplete input using Code LLMs, which have been integrated into commercial applications [19, 2].
+ShadowCode achieves an ASR of up to 93.3% on CodeGeeX and 90.4% on Copilot using perturbations generated by open-source Code LLMs, confirming its serious threat in the real world.
+In summary, this paper makes three main contributions: (1) We reveal the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduce a new attack paradigm, i.e., external prompt injection.
+(2) We propose a simple yet effective method (i.e., ShadowCode).
+ShadowCode can generate concise induced perturbations with effectiveness in an uncontrolled context environment.
+(3) We conduct extensive experiments to evaluate ShadowCode under 31 representative threat cases on 3 representative open-source Code LLMs and 2 commercial Code LLM-integrated applications, demonstrating its effectiveness, stealthiness, and resistance to potential defenses.
+2 Background and Related Work
+2.1 Code LLMs for Code Completion
+Code and its comments serve as a valuable corpus for deep learning models, facilitating various code-related tasks, including code translation [29] and code understanding [30].
+The emergence of LLMs provides new alternatives to coding tasks.
+Leveraging the extensive knowledge and reasoning capabilities of LLMs, researchers have developed specialized models, e.g., Codex [3], CodeLlama [31], CodeGemma [4] and CodeGeex [5].
+These Code LLMs can provide detailed explanations regarding the intrinsic logic of code and efficiently generate a wide range of functionally complex code [32, 33].
+This paper focuses on code completion [7, 6].
+This is a paradigmatic task involving the generation of new code from incomplete input using Code LLMs, which have been integrated into commercial applications [1, 2].
 However, this development paradigm may encounter some potential threats.
-On the one hand, Code LLMs can be easily influenced by the code context, which alters the result of code generation.
+On the one hand, Code LLMs are susceptible to influence from the code context, which can alter the generated code.
 On the other hand, attackers can easily control the context of code generation, such as by inserting external code when the user copies code, installs packages, or builds knowledge bases, thus affecting the decisions made by Code LLMs.
-These threats have already occurred in the real world [7], highlighting the necessity of its security research.
-II-B Backdoor and Adversarial Attacks
-Backdoor attacks target the training process of DNNs [20, 31, 58].
-These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [26, 30, 21].
+Related events have already occurred in the real world [34], highlighting its realistic threat.
+2.2 Backdoor and Adversarial Attacks
+Backdoor attacks target the training process of DNNs [35, 36].
+These attacks implant malicious backdoors that can be activated by an attacker-specified trigger into a model during training, leading the model to misbehave when the trigger is present [37, 38].
 In the code domain, there are generally two strategies to achieve this, including dataset poisoning [10, 12] and model poisoning [11].
@@
 Recently, a few works also exploited backdoor attacks for positive purposes [39, 40, 41], which are out of our scope.
-Adversarial attacks target the model’s inference process [11, 43, 25].
-Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [59, 28, 47].
-In the code domain, existing adversarial attacks [12, 62, 24] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs.
+Adversarial attacks target the model’s inference process [42, 43].
+Adversarial attacks introduce perturbations to the input data, leading the model to make incorrect output [44, 45].
+In the code domain, existing adversarial attacks [46, 14] aim to reduce the quality of generated code by crafting adversarial examples based on the (approximated) gradients regarding the model’s inputs and outputs.
 For example, Jha et al.[14] slightly changed the original code (e.g., function and variable names) to lower the performance of Code LLMs measured by CodeBLEU score [47].
@@
 Accordingly, adversarial attacks are difficult to pose a serious threat to Code LLMs in practice.
-II-C Prompt Injection and Indirect Prompt Injection Attacks
-Prompt injection (PI) attacks, are a special form of attack targeting LLMs [15, 36, 63].
-A malicious LLM user aims to bypass the safety guardrails of the LLM and forces it to perform undesired responses, e.g., details of creating bomb.
-In the code domain, existing research [39, 37, 17] focuses on code-oriented attack objectives, such as using LLMs to create malware.
-In general, these attacks target the developers of models (to make their models unsafe) instead of users, which is out of the scope of this paper.
-Indirect prompt injection (IPI) attack attempts to corrupt the prompt submitted by the user to the LLM and uses the LLM as an intermediary to output malicious responses, e.g., recommending phishing websites [22].
-This attack simultaneously threatens both LLM-integrated applications (by reducing usability) and victim users (by receiving malicious responses).
-In the code domain, the only existing work [22] makes some preliminary attempts: it presents two simple demos without the implementation of malicious functionalities; it does not consider stealthiness and inserts a large amount of readable comments containing the malicious intent.
-Arguably, the existing IPI attack does not pose a realistic threat to Code LLMs.
-III Problem Statement
-III-A Code Completion Task
-Code completion is an essential task for Code LLMs.
-The input to the models may include code snippets, comments, pseudocode, and natural language requests.
-The output should consist of executable code blocks that meet the requirements specified by the input.
+2.3 Prompt Injection Attacks
+Prompt injection is a type of input manipulation attack against LLMs [48, 18], where attackers embed carefully crafted instructions in the dialogue prompt or context, causing the model to violate established system policies or developer intentions, leading to undesired responses (e.g., illeagal content [17], information leakage [49] and unauthorized operations [18]).
+In the code domain, existing research typically focus on the attacks conducted by malicious users, such as exploiting the Code LLMs to create malware [16].
+These attacks aim for “jailbreaking” and are not concerned with the threat prompt injection poses to normal users of Code LLM-integrated application.
+Meanwhile, we notice that there are some methods sharing a similar attack scenario with ours, named indirect prompt injection (IPI) [18, 50], where attackers strategically inject the malicious instructions into the data likely to be retrieved at LLMs’ inference time.
+However, these attacks typically involve manually designing injection instructions (e.g., demanding disregard for the system prompt) and rely on a covert information space (e.g., the Markdown of a Wikipedia page) without considering perturbation stealthiness.
+This leads to two problems in the real world: (1) Manually designed injection instructions lack flexibility and sometimes even effectiveness.
+(2) It’s often difficult to find a covert information space in many scenarios, e.g., in code files.
+Currently, in the code domain, exsiting IPI has only provided some preliminary showcases [18].
+As shown in Figure 12 in Appendix A, researchers used 23 lines of comment to inject data deletion instructions (which could be achieved with just 10 tokens using ShadowCode), indicating a lack of substantial real-world threat.
+Figure 2: The overall pipeline of our ShadowCode for (automatic) external prompt injection attack against Code LLMs.
+ShadowCode is executed in two stages: (1) Design Optimization Objectives: given the attacker’s malicious objective, ShadowCode converts it into a specific output tuple (position code, target code).
+Subsequently, ShadowCode expands this output tuple to simulate a code context of the victim user’s project, based on which to design the loss function of perturbation optimization.
+(2) Perturbation Optimization: after defining the loss function, we use a surrogate Code LLM to calculate the loss and employ a greedy search to optimize the induced perturbation token by token.
+In particular, we enhance the optimization with forward reasoning enhancement (FRE) and keyword-based perturbation design (KPD) to improve the perturbation’s effectiveness and stealthiness, respectively.
+After obtaining the induced perturbation through ShadowCode, the attacker inserts it into a malicious payload, which can then be strategically introduced into the code context via various channels (e.g., direct copy, dependencies, or RAG).
+Subsequently, this perturbation within the code context will induce Code LLMs to generate specific code at an attacker-specified position, thereby achieving the malicious objective.
+3 Methodology
+3.1 Formulation of External Prompt Injection Attacks
+The Main Pipeline of Code Completion Task.
+Code completion is an essential task for Code LLMs, whose input may include code snippets, comments, pseudocode and user-specified natural language requests.
+Its output should consist of executable code blocks that meet user-specified requests.
 In this paper, we consider a general task format.
-Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},x_{3},...,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements.
+Specifically, the input code block is defined as $\bm{X}=\{x_{1},x_{2},\ldots,x_{m}\}$, where $x_{i}\in\bm{X}$ can be any token, and $\bm{X}$ as a whole adheres to the code format requirements.
 For instance, natural language requests can only be included in the code comments.
-The output is the completed code block $\bm{Y}=\{x_{1},x_{2},x_{3},...,x_{m},x_{m+1},x_{m+2},...,x_{n}\}$, which should be logically executable and functionally compliant with the requirements.
+The output is the completed code block $\bm{Y}=\{y_{1},y_{2},\ldots y_{n}\}$, which should be logically executable and functionally compliant with the requirements.
 We denote a Code LLM with weights $\bm{\theta}$ as $\Theta$, and the code completion task is simplified as $\bm{Y}=\Theta(\bm{X})$.
-III-B Target-specific Attack
-In the code domain, target-specific attack refers to the scenario where the attacker makes the Code LLM generate unsafe code for specific targets.
-Generally, the attacker’s targets can be described as a tuple of two continuous token sequences $(\bm{Y_{T}},\bm{T})$, where $\bm{Y_{T}}=\{x^{t}_{1},x^{t}_{2},...,x^{t}_{j}\}$ is a subsequence within $\bm{Y}$, and $\bm{T}=\{t_{1},t_{2},t_{3},...,t_{k}\}$ is what the attacker aims to insert after $\bm{Y_{T}}$ to create malicious functionality.
-To achieve the targets, the attacker needs to design a perturbation $\bm{P}$, represented as $\bm{P}=F(\bm{Y_{T}},\bm{T})$, and insert it into the input.
-The function $F$ can be implemented as manually creating the perturbation or by an optimization-based method as proposed in this paper.
-It is worth noting that, unlike backdoor attacks, the connection between $\bm{Y_{T}}$ and $\bm{T}$ is not built during the training process.
-To conclude, for a successful target-specific attack, the model output received by the victim $\bm{Y_{v}}$ will be:
-$\begin{split}Y_{v}&=\Theta(\bm{X}\oplus\bm{P})\\ &=\{x_{1},x_{2},...,x^{t}_{1},x^{t}_{2},...,x^{t}_{j},t_{1},t_{2},...,t_{k},...,x_{n}\},\end{split}$ (1)
-where $\oplus$ means contacting the two token sequences.
-Figure 2: The overall pipeline of our TPIA attack.
-It is executed with two attack steps.
-(1) Objective Transfer.
-The attacker constructs the optimization task based on the malicious objective and generates a concise non-functional perturbation (instead of functional code used in supply chain attacks), containing instructions to achieve the constructed objective.
-(2) Perturbation Injection.
-The malicious code containing the malicious perturbation will be submitted to the open-source platforms (e.g., GitHub [3] and StackOverflow [5]) and introduced to the victim’s code context via various paths, such as direct copy and installing packages.
-When the victim requests the Code LLM to perform code completion based on his code context, the perturbation will induce the model to respond to the malicious instructions at specific locations.
-III-C Threat Model
-Attack Scenarios.
-We consider common scenarios where users introduce external code by copying or installing pack- ages [22, 41].
-The attacker generates a non-functional perturbation to circumvent detection methods (e.g., static analysis [33] and sandbox [9]) and inserts it into victim’s code context via the external code.
-When the user requests the Code LLM to perform code completion, the perturbation will induce the model to generate an unsafe output at the specified location.
-Besides introducing code to the victim’s local project, attackers can also upload their codes containing perturbation to popular open-source platforms (e.g., GitHub), whose code will be automatically downloaded as a knowledge base to enhance the performance of Code LLMs [13] via retrieval-augmented generation [27].
-Perturbation in the knowledge base can also appear in the context of the Code LLM’s prompts.
-Attacker’s Capabilities.
-We assume that the attacker can obtain the parameters of an assistive Code LLM to generate malicious perturbations.
-This assumption is consistent with many adversarial methods [63, 50, 24].
-Specifically, the attacker ideally obtains the open-source model corresponding to the target application, or uses the most effective assistive model in gray-box and black-box scenarios.
-On the other hand, perturbations can be detected and removed.
-Therefore, the attacker needs to employ non-functional perturbations rather than the functional code commonly used in supply chain attacks [41], since the latter can be easily detected by existing methods (e.g., static analysis [33] and sandboxing [9]).
-Attacker’s Goals.
-The attacker has a specific malicious objective, which can be represented as a tuple (target position, target code).
-This tuple follows the definition of target-specific attacks in Section III-B.
-Specifically, the target position can be represented as $\bm{Y_{T}}$, and the target code can be represented as $\bm{T}$.
-The attacker attempts to insert a piece of designed malicious code $\bm{T}$ after a segment of the victim’s normal code $\bm{Y_{T}}$, and cause the victim to receive abnormal output code $\bm{Y_{v}}$.
-To achieve this goal, the attacker applies the adversarial methods to compute a perturbation $\bm{P}$ from the Code LLM’s parameters $\bm{\theta}$, which is most likely to induce the Code LLM to complete $\bm{T}$ after $\bm{Y_{T}}$.
-It can be formulated as below:
-$\begin{array}[]{c}\bm{P}=\text{Adv}(\bm{\theta},\bm{Y_{T}},\bm{T}).\end{array}$ (2)
-Note that Equation (2) can be regarded as the ‘adversarial attack version’ of $F(\bm{Y_{T}},\bm{T})$ defined in Section III-B and the generated perturbation $\bm{P}$ should satisfy Equation (1).
-In particular, the attacker can achieve these goals by inserting code snippets into specified positions, without needing to oppose the Code LLMs’ safety alignment.
-For example, when setting the target position to DES.new( and the target code snippet to key,DES.MODE_ECB), the generated perturbation can induce Code LLMs to use the easily cracked ECB mode for encryption; When setting the target position to TRG() (a shorthand for Target which can be replaced with any function name), the generated perturbation can induce Code LLMs to insert designed code snippet into the TRG function or after using it.
-IV Methodology
-IV-A Overview
-We hereby briefly introduce our target-specific prompt injection attack (TPIA).
-In general, TPIA consists of two main stages: task construction and malicious perturbation generation, as shown in Figure 2.
-Specifically, in the first stage, we construct the input and target output based on the attacker-specified objective (Section IV-B).
-In the second stage, we generate the malicious non-functional perturbation to achieve the constructed objective (Section IV-C).
-The perturbation generation process in the second stage involves two optimizations: (1) forward reasoning enhancement aligns the optimization goals of adversarial computation with Code LLMs’ sampling reasoning process (Section IV-D); (2) keyword-based perturbation design reduces the token usage of the perturbation to enhance the stealthiness (Section IV-E).
-We describe their technical details in the following parts.
-IV-B Task Construction
-In the first stage, TPIA constructs specific code completion tasks based on the malicious objectives.
-The constructed tasks simulate the victim’s local code environment and are used to optimize the malicious perturbation, which is closely related to the final performance of TPIA.
-Each task can be divided into two parts: the input containing the target position, and the expected output containing the target code.
-This task is to be constructed so that our attack satisfies the following three properties to ensure effectiveness.
-Property 1: Feasible Malicious Intent.
-The code environment configured by the user must meet the basic requirements for the malicious objective, including the necessary packages and methods declarations.
-Therefore, we introduce the conditional code at the beginning of the input code.
-Property 2: Accurate Code LLMs’ Response.
-The response of the Code LLMs should be the target position.
-As stated in Section III, the target code should be inserted after the target position to achieve the malicious objective.
-Hence, we place the position code at the end of the input, maintaining its continuity with the target code.
-Property 3: Robust Malicious Perturbation.
-The attacker cannot anticipate user’s contextual code in advance, which is the inevitable noise for TPIA.
-Therefore, we design the context code and add irrelevant noise to enhance the robustness of the malicious perturbation.
-The context code is placed between the position code and conditional code.
-To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode, i.e., making the Code LLM insert Target Code (‘‘DES.MODE_ECB’’) after “cipher = DES.new(key,”.
-To achieve this, TPIA constructs an output tuple as (“cipher = DES.new(key,” , “DES.MODE_ECB)”.
-The other three task components are automatically generated with the following stages.
-Firstly, we declare the Crypto series methods that enable the encryption code to run normally in the conditional code.
-Secondly, we add “cipher = DES.new(key,” to the last line of the position code to ensure the continuity with the target output “DES.MODE_ECB)”.
-In this example we also consider the necessary parameter “key” for the encryption method, as well as the disturbance parameter “iv”, which is required by encryption methods like “CBC” but not needed by “ECB”.
-Finally, TPIA constructs the code unrelated to the task and uses it as the noise-like context code.
-Particularly, TPIA constructs context codes of multiple aspects for the same task to collaboratively enhance the robustness.
-Figure 3: The example of task construction that induces Code LLM to use encryption mode with low security.
-In the entire demo file, TPIA simulates the parts related to the malicious objective and constructs the tasks accordingly.
-IV-C Malicious Perturbation Generation
-In the second stage, TPIA optimizes the perturbation through an adversarial gradient computation method such that it can most likely induce the Code LLMs to generate the target malicious code.
-Utilizing the task code constructed in the first stage and the probabilities output by the target model, TPIA calculates the gradient of each token in the perturbation at the embedding layer.
-Through a greedy search approach, the optimal perturbation is derived.
-The optimization process is illustrated in Algorithm 1, with detailed descriptions below.
-Algorithm 1 Perturbation Tokens Optimization
-1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{T}$, initial perturbation $\bm{P}$ with adversarial parts $x_{\text{adv},1:l}$, iterations $k$
-2:  Vocab set $V$, $\mathbf{e}\in V$
+The Formulation of EPI.
+In general, there are three parties, including a Code LLM-integrated application, a user, and an attacker, involved in our EPI.
+The user employs the Code LLM-integrated application to complete the code project.
+The Code LLM-integrated application constructs the input $\bm{X}$ for the integrated Code LLM based on the code project and get the output.
+The attacker generates a concise induced perturbation $\bm{P}$ and introduces it into the input $\bm{X}$.
+The Code LLM, affected by $P$, will generate code that is functionally close to the attacker’s malicious objective $O$, i.e.,
+$\mathcal{D}(O,\mathcal{F}(\Theta(\bm{X}\leftarrow\bm{P})))<\epsilon,$ (1)
+where $\mathcal{F}(\cdot)$ denotes the functionality of the code, $\bm{X}\leftarrow\bm{P}$ means injecting $\bm{P}$ into $\bm{X}$.
+$\mathcal{D}(\cdot,\cdot)$ is a function that measures functional differences.
+When its value is less than $\epsilon$, it indicates functional equivalence between inputs.
+Attacker’s Capacities and Attack Scenarios.
+We assume that the attacker can obtain the parameters of an surrogate Code LLM to conduct attacks.
+This assumption is consistent with many adversarial methods [26, 51].
+Specifically, we consider three main attack settings: white-box, gray-box, and black-box settings.
+The former assumes that the attacker knows the specific parameters of Code LLM, which generally corresponds to scenarios where the model is an open-sourced model.
+The latter two assume that the attacker only knows the model series (rather than the specific version) and knows nothing about it, respectively, which generally corresponds to scenarios of using commercial Code LLM-intergrated applications.
+In our attack scenarios, the attackers introduce external code to the victim user’s code project.
+We argue that it can be easily and stealthily achieved in the real world.
+For example, introducing external code by copying or installing packages is commonly done during the programming process [23, 18].
+Meanwhile, external code from open-source platforms (e.g., GitHub) is usually downloaded as a knowledge base via retrieval-augmented generation [22], which can also be used as input to improve the performance of Code LLMs [21].
+In particular, to avoid detection by existing methods (e.g., static analysis [24] and sandboxing [25]), external code should not immediately contain malicious functionalities, i.e., attackers should use non-functional code as induced perturbation.
+3.2 Overview
+In this section we briefly introduce our greedy perturbation optimization method (dubbed ‘ShadowCode’) for automatic and effective external prompt injection.
+As illustrated in Figure 2, ShadowCode operates in two sequential stages: (1) design optimization objectives and (2) perturbation optimization.
+In the first stage, we design the perturbation optimization objectives for induced perturbation based on the attacker’s malicious objective.
+In the second stage, we employ the greedy search to optimize the induced perturbation.
+3.3 Design Optimization Objeciteves
+In this stage, we first describe how to convert diverse malicious objectives to structured output tuples.
+Then we expand the output tuple to a full code contextual simulation.
+Finally, we design a loss function that leverages the simulation to guide the search for the optimal perturbation.
+Output Tuple.
+We hereby describe how to convert the malicious objective to an output tuple.
+Let the normal code completed by a Code LLM is denoted by $\bm{Y}$, the implementation of malicious purposes is maliciously modifying $\bm{Y}$.
+Such modifications can be divided into three types: deletion, addition, and alteration, all of which can be represented by an output tuple (position code $\bm{Y}_{P}$, target code $\bm{Y}_{T}$).
+Specifically, for deletion, the attacker wants to delete the code segment $[y_{i},y_{i+1},\ldots,y_{i+k}]$ in $\bm{Y}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$ and $\bm{Y}_{T}=[y_{i+k+1},y_{i+k+2},\ldots,y_{t}]$; For addition, the attacker wants to add the code segment $\bm{T}$ into the position between $y_{i}$ and $y_{i+1}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$ and $\bm{Y}_{T}=[\bm{T},y_{i+1},y_{i+2},\ldots,y_{t}]$; For alteration, the attacker wants to change the code segment $[y_{i},y_{i+1},\ldots,y_{i+k}]$ to $\bm{T}$, and we have $\bm{Y}_{P}=[y_{p},y_{p+1},\ldots,y_{i-1}]$, $\bm{Y}_{T}=[\bm{T},y_{i+k+1},y_{i+k+2},\ldots,y_{t}]$.
+Code Contextual Simulation.
+Optimizing the induced perturbation based solely on the output tuple ignores the surrounding code context.
+To ensure the effectiveness of ShadowCode in uncontrolled context environment, we expand the output tuple to simulate a code context of the victim user’s porject.
+In general, although the contextual environment of the victim user’s project is usually uncontrollable for the attacker, there are still some fundamental logics to follow for code projects and code generation processes, which can be exploited by ShadowCode to simulate a realistic environment.
+Accordingly, our code context simulation is constructed based on the following three logics:
+Logic 1: Necessary Condition Argument.
+The code environment configured by the user must meet the basic requirements for the malicious objective, such as the necessary packages and method declarations.
+Therefore, we introduce this kind of code as conditional code at the beginning of the whole simulated code.
+Logic 2: Fixed Generation Order.
+Code generation usually proceeds in a forward textual order.
+Accordingly, the target code should be inserted immediately after the position code to achieve the malicious objective and the subsequent content of the target code is not strongly relevant to the objective.
+Thus, we connect the position code and target code and place them at the end of the simulated code.
+Logic 3: Noise in Context.
+Since realistic projects usually consist of code with different functions, the code context is likely to contain noise code not entirely related to malicious purposes, which may have a potential impact on the effectiveness of the final induced perturbation.
+As such, we design the noise code, adding irrelevant noise to enhance the robustness of the induced perturbation.
+Based on three logics above, we can construct the code contextual simulation for different malicious objectives.
+To better illustrate our strategy, we provide a straightforward example in Figure 3, where the malicious objective is to make the victim use a weak encryption mode.
+The task components are automatically generated with the following stages.
+First, the position code should contain a fundamental encryption situation: the encryption key “key”, the text to be encrypted “iv” and the first half of the encryption instruction “cipher = DES.new(key,”.
+These variables and functions are necessary for the encryption methods, e.g., “CBC”.
+To achieve the malicious objective, the complete code instruction should call the weaker encryption method, “ECB”, thus the target code is “DES.MODE_ECB)”.
+After obtaining the output tuple, ShadowCode then declares the Crypto series methods that enable the encryption code to run normally in the conditional code.
+Finally, ShadowCode constructs the code that is not related to the task and uses it as the noise code.
+The final code contextual simulation is constructed from multiple aspects for the same task to collaboratively ensure the effectiveness.
+Figure 3: The example of code contextual simulation, where the attacker aims to induce the Code LLM to use encryption mode with low security (i.e., ECB encryption).
+The simulation consists of four code segments: conditional code, noise code, position code, and target code, to simulate the context environment most relevant to the malicious objective.
+Loss Function of Perturbation Optimization.
+After obtaining the code contextual simulation, a loss function can be formulated for optimizing induced perturbations.
+To achieve the malicious objective, the optimization aims to identify the perturbation that maximizes the probability of realizing the simulated successful attack.
+Specifically, the simulated contextual code is first divided into input and output components.
+The input to the Code LLM comprises conditional code, noise code, and position code, simulating a code completion task that can be exploited to achieve the malicious objective.
+This entire input is referred to as task code.
+The output, termed target code, serves as the label for the perturbation optimization task.
+Subsequently, with $\bm{Y}_{T}=[t_{1},t_{2},\ldots t_{k}]$, the loss function is defined as:
+$\mathcal{L}=$ $\sum_{i=1}^{|\bm{Y}_{T}|}\log(1-\mathbb{P}(t_{i}|\bm{T}_{s}\leftarrow\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (2)
+where $\bm{T}_{s}$ represents the task code, the probability $\mathbb{P}$ is inferred and calculated using the surrogate model weights $\bm{\theta}$.
+This function aims to maximize the likelihood of generating the target code given the task code as input.
+When the Code LLM receives the perturbation, it is induced to generate the target code if the context is appropriate (i.e., similar to the simulation in code functionality), thereby achieving the attacker’s malicious objectives.
+3.4 Perturbation Optimization
+In this stage, we first introduce two enhancement modules for the optimization process: forward reasoning enhancement and keyword-based perturbation design.
+Then, we will describe the whole algorithm for induced perturbation optimization.
+Forward Reasoning Enhancement.
+Directly employing the loss function defined in Equation (2) may lead to distortions during the autoregressive generation process.
+As shown in Figure 4, since token prediction results are based on previous sampling outcomes, the original loss function may deviate from actual conditions due to unexpected predictions of the first few tokens.
+For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context (i.e., immediately after calling a demo function “TRG”).
+If the induced Code LLM fails to generate the first token “os”, the probabilities of generating subsequent tokens will diverge significantly.
+Figure 4: Distortion during the autoregressive generation process.
+The attacker aims to insert a malicious deletion command after the position code, whereas the Code LLMs would typically generate the result printing command.
+During the perturbation optimization process, the generation probability of the first token in the target code, i.e., “os”, is not adequately optimized.
+Consequently, the Code LLM may not be successfully induced upon encountering the position code and generate an undesired token “print”.
+Subsequently, the failure to generate the first token compromises the successful generation of subsequent tokens.
+To alleviate this problem, a forward reasoning enhancement module is introduced.
+Motivated by the aforementioned discussion that subsequent tokens are influenced by preceding tokens in an autoregressive generation manner, earlier tokens hold greater importance for the overall generation process of the target code.
+Therefore, the importance of the initial tokens for the loss value is increased by assigning them higher weights, thereby enabling the optimization objective to adapt to the autoregressive mechanism.
+This increase in weight can be achieved through adding an additional loss.
+Its specific formulation is shown as follows:
+$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-\mathbb{P}(t_{i}|(\bm{T_{s}}\leftarrow\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (3)
+$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
+where $\mathcal{L}_{e}$ is the additional loss, $r$ is its coefficient, and $h$ is the number of enhanced tokens.
+Figure 5: Examples of different induced perturbations.
+The attacker aims to induce the Code LLM to generate a data deletion command (“os.system("rm -rf /")”) immediately after a specified function call (“TRG”) via induced perturbation.
+Manually designed induced perturbations lack stealthiness due to their obvious malicious intent, a flaw that a simple greedy search-based replacement alone cannot rectify.
+Induced perturbations optimized with random initialization necessitate extensive token usage to ensure effectiveness.
+Our method combines optimized induced perturbations with a prefix (keyword), thereby achieving both effectiveness and stealthiness with limited tokens.
+Keyword-based Perturbation Design.
+This module aims to significantly reduce the token count required for induced perturbations by offloading information into prefix tokens, which enhances attack stealth.
+Arguably, minimizing token usage and obscuring malicious characteristics are critical to maintaining attack stealth, thereby avoiding detection by victim users or RAG’s developers.
+Generally, perturbations optimized via greedy search typically exhibit less obvious malicious characteristics compared to manually designed ones.
+However, these perturbations struggle to accurately convey malicious instructions as effectively as manually designed perturbations, resulting in higher token usage.
+A seemingly straightforward solution, which involves optimizing based on manually designed perturbations by replacing obvious malicious characteristics with equivalent tokens, proves difficult in practice.
+As shown in Figure 5, the tokens constituting these obvious malicious characteristics (i.e., convey dangerous data deletion instructions) are the most critical and challenging to replace via greedy search.
+To alleviate these limitations, we propose a simple yet effective enhancement method, i.e., keyword-based perturbation design.
+Intuitively, the purpose of the induced perturbation is to fulfill the attacker’s malicious objective.
+Therefore, the output tuple, which expresses malicious objective, will contain the necessary information for the induced perturbation.
+Motivated by this understanding, we leverage tokens from the output tuple to construct a prefix for the perturbation, termed a keyword, which is then injected into the victim’s code context alongside the perturbation.
+During keyword construction, we use the loss function in Equation (3) to greedily select the optimal token combination.
+In particular, we utilize only a very small number of tokens (typically one or two) for keyword construction to ensure no overt malicious intent is present in the injected content.
+As shown in Figure 5, the keyword “TR” exhibits no overt malicious intent but can enhance the perturbation’s effectiveness compared to directly optimized counterparts.
+Optimization Algorithm.
+We hereby introduce the specific optimization process of ShadowCode for induced perturbations and their corresponding keywords.
+Specifically, we employ greedy gradient search to optimize an initialized perturbation $\bm{P}=[p_{1},p_{2},\ldots p_{l}]$, where $l$ denotes the token count.
+As shown in Algorithm 1, our ShadowCode method optimizes the perturbation token by token in each iteration (line 4).
+We utilize the loss function in Equation (3) for greedy gradient search (line 5), i.e., the optimization objective for the $i$-th token $p_{i}$ is:
+$\begin{array}[]{c}p_{i}=\mathop{\arg\min}\limits_{p_{i}}\mathcal{L}.\end{array}$ (4)
+To achieve this, ShadowCode propagates the loss value to the vocabulary embedding layer.
+Based on the gradient at token $p_{i}$, it greedily searches for the suitable token $e^{\prime}_{i}$ within the Code LLM’s vocabulary set $V$ to replace $p_{i}$, as shown in the following equation:
+$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-p_{i}\right)^{T}\nabla_{p_{i}}\mathcal{L},$ (5)
+where $\nabla_{p_{i}}\mathcal{L}$ represents the gradient of the task loss.
+ShadowCode enhances this token replacement strategy by employing beam search, which considers the Top-$k$ token candidates (Line 6-7).
+ShadowCode then replaces the original token with each candidate to form a new $\bm{P^{\prime}}$ (Line 8-9), and computes a new loss value $\mathcal{L}^{\prime}$: if $\mathcal{L}^{\prime}$ is better than $\mathcal{L}$, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13).
+Finally, if $\bm{P}$ is not optimized within an entire iteration, ShadowCode considers it optimal and terminates the optimization process (Line 16).
+Algorithm 1 Induced Perturbation Optimization
+1:  Input: Task Code $\bm{T_{s}}$, target code $\bm{Y}_{T}$, initial perturbation $\bm{P}$ with tokens $p_{1:l}$, beam search size $k$
+2:  Vocab set $V$, weights $\bm{\theta}$ of the surrogate Code LLM
 3:  repeat
 4:     for $i\in[0,\ldots,l]$ do
-5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P},\bm{\theta})$
-6:        $\mathcal{G}=\nabla_{x_{\text{adv},i}}\mathcal{L}$
-7:        $\mathbf{x_{\text{candidates}}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
+5:        $\mathcal{L}=\text{get\_loss}(\bm{T_{s}},\bm{Y}_{T},\bm{P},\bm{\theta})$
+6:        $\mathcal{G}=\nabla_{p_{i}}\mathcal{L}$
+7:        $\mathcal{P}_{\text{candidates}}=\text{Top-}k_{e^{\prime}_{i}}\text{min}({e^{\prime}_{i}}^{T}\mathcal{G}),e^{\prime}_{i}\in V$
 8:        for $j\in[0,\ldots,k]$ do
-9:           $\bm{P}^{\prime}:x_{\text{adv},i}\to x_{\text{candidates}}^{j}$
-10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{T},\bm{P^{\prime}},\bm{\theta})$
+9:           $\bm{P}^{\prime}:\text{replace }p_{i}\text{ to }\mathcal{P}_{\text{candidates}}^{j}$
+10:           $\mathcal{L}^{\prime}=\text{get\_loss}(\bm{T_{s}},\bm{Y}_{T},\bm{P^{\prime}},\bm{\theta})$
 11:           if $\mathcal{L}^{\prime}<\mathcal{L}$ then
@@
 15:     end for
-16:  until perturbation did not change in the last iteration
+16:  until $\bm{P}$ does not change in the whole iteration
 17:  Output: Optimized perturbation $\bm{P}$
-Firstly, TPIA sets the adversarial token count $l$ and randomly initializes the perturbation.
-Considering that our enhancement method adds static tokens to the perturbation (Section IV-E), here we denote the adversarial part of the perturbation as $x_{\text{adv}_{1:l}}$.
-After determining the input and output of the code completion task, TPIA starts to optimize the $x_{\text{adv}_{1:l}}$ token by token in each iter (Line 4).
-We design a loss function based on the probability of the Code LLM for generating the target output (Line 5), which can be formulated as follows:
-$\mathcal{L}=$ $\sum_{i=1}^{|\bm{T}|}\log(1-p(t_{i}|\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{i-1})),$ (3)
-where $\bm{T_{s}}$ denotes the task code.
-We only consider the prediction probability of each token in the target code when designing the loss function since the success of TPIA is solely related to whether the target code is correctly generated.
-Besides, the optimization objective for the $i$-th adversarial token $x_{\text{adv}_{i}}$ can be denoted as:
-$\begin{array}[]{c}x_{\text{adv}_{i}}=\mathop{\arg\min}\limits_{x_{\text{adv}_{i}}}\mathcal{L}.\end{array}$ (4)
-To achieve this goal, TPIA propagates the loss value to the vocabulary embedding layer.
-Based on the gradient at token $x_{\text{adv}_{i}}$, it greedily searches for the suitable token $e^{\prime}_{i}$ in the vocabulary set of the Code LLM $V$ to replace $x_{\text{adv}_{i}}$, as shown in the following equation:
-$e^{\prime}_{i}=\underset{e^{\prime}_{i}\in V}{\arg\min}\left(e^{\prime}_{i}-x_{\text{adv}_{i}}\right)^{T}\nabla_{x_{\text{adv}_{i}}}\mathcal{L},$ (5)
-where $\nabla_{x_{\text{adv}_{i}}}Loss$ is the gradient of the task loss.
-TPIA uses beam search to enhance this token replacement strategy by considering the Top-$k$ token candidates (Line 6-7).
-This has been proven to be highly effective in existing NLP tasks [63, 50].
-Afterwards, TPIA uses each token candidate to replace the original one to form a new $\bm{P^{\prime}}$ (Line 8-9).
-It then computes the new loss value using $\bm{P^{\prime}}$ and compares it with the original one: if the new value is better, $\bm{P}$ is replaced with $\bm{P^{\prime}}$ (Line 10-13).
-Finally, if $\bm{P}$ is not optimized within an entire cycle from 0 to $l$, TPIA considers it as optimal and terminates the optimization process (Line 16).
-In particular, the design of our malicious perturbation generation is inspired by existing methods, such as universal trigger [50] and GCG [63].
-We have already made some special designs, e.g., using Positon Code to locate the specific malicious target, based on characteristics of code completion and indirect prompt injection (instead of prompt injection).
-However, its effect is still very limited due to the difficulties of targeted malicious tasks.
-For example, as shown in Figure 9, its attack success rate (0-key, no FRE) is 0%.
-To address this, we propose Forward Reasoning Enhancement and Keyword-based Perturbation Design to further improve the perturbation generation process of our TPIA.
-Figure 4: Loss function distortion during the forward reasoning process.
-When calculating the loss value using Equation 3, it optimizes the highest probability for generating the target token sequence.
-However, if the generation of the first target token fails, it can cause the loss function to become entirely inconsistent with the actual situation, making the perturbation ineffective in practical scenarios.
-IV-D Forward Reasoning Enhancement
-This optimization allows the optimization direction of adversarial computation to be closer to the actual reasoning process of Code LLMs (i.e., forward token sampling with randomness), thus improving the attack’s effectiveness and robustness.
-As shown in Figure 4, due to the operational mode of LLMs, where the final result is based on former sampling results, the original optimization function may deviate from actual conditions in certain situations.
-For example, the data deletion malicious instruction “os.system("rm -rf /")”, although brief, is difficult to integrate coherently with the surrounding context.
-This phenomenon is highly related to the optimization design in Equation (3), where TPIA calculates and sums the generation probabilities of individual tokens based on predetermined conditions.
-This indicates that a deviation occurring during the sampling of preceding tokens can distort the probability calculations for subsequent tokens.
-To alleviate this problem, we propose a forward reasoning enhancement technique.
-Our intuition is that the former tokens of the target code should be assigned more weights to enhance the forward reasoning.
-Considering the computational overhead, we assign higher weights to the first $h$ tokens to mitigate the gradient calculation direction error.
-Its specific formulation is shown as follows:
-$\mathcal{L}_{e}=\frac{1}{h}\sum_{i=1}^{h}\sum_{j=1}^{i}$ $\log(1-p(t_{i}|(\bm{T_{s}}\oplus\bm{P},t_{1},t_{2},\ldots,t_{j-1}))),$ (6)
-$\mathcal{L}=(\mathcal{L}+r\mathcal{L}_{e})/(r+1),$
-where $\mathcal{L}_{e}$ represents the additional reasoning-based enhancement applied to the original loss function, and $h$ is a hyperparameter set to either 1 or 2, which significantly improve the performance of attack in practice.
-Figure 5: Examples of the keyword-based perturbation and other perturbations.
-The directly generated perturbation faces challenges in being effective while the manually-crafted perturbation lacks the stealthiness advantage.
-Our keyword-based perturbation can balance both effectiveness and stealthiness.
-IV-E Keyword-based Perturbation Design
-This optimization offloads insensitive parts of the malicious objective into static tokens, significantly reducing the total number of tokens required for a successful attack.
-From an empirical perspective, using fewer tokens and less obvious malicious characteristics will enhance the stealthiness of the attack.
-Compared to the manually-crafted perturbations, the illegibility of our malicious perturbations can prevent the victim from noticing their malicious intent.
-However, accurately conveying instructions using illegible tokens remains a significant challenge.
-A straightforward solution is to use a manually-designed perturbation as initialization, but this approach has two issues.
-First, the malicious perturbation cannot gain any advantage in terms of the token count.
-Second, this leads to ineffective optimization efforts.
-As shown in Figure 5, the manually-designed perturbation fails to not only reduce the token usage but also optimize out the core malicious features, leading to additional optimization rounds.
-To alleviate these limitations, we propose another simple yet effective enhancement method, i.e., keyword-based perturbation design.
-Our idea is to transfer part of the information to static tokens, thereby reducing the information load on illegible tokens.
-Specifically, before each round of perturbation optimization in Algorithm 1, we use grid search to select the n tokens with the best performance as keywords.
-With the optimization step $s$, the equation can be formulated as follows:
-$\bm{P}_{key}^{s+1}=\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}_{adv}^{s},\bm{\theta}),$ (7)
-$\bm{P}_{adv}^{s+1}=\min_{\bm{P}_{adv}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P}_{adv},\bm{\theta}),$
-where $\bm{P}_{key}$ and $\bm{P}_{adv}$ represents keywords part and adversarial part in perturbation $\bm{P}$ and the loss function is the same as Equation 6.
-To make the keywords relevant to the malicious objective, we select the tokens for the keywords from the target code and the position code, i.e., for token $p_{ki}$ in $\bm{P}_{key}$, there is $p_{ki}\in\{\text{\emph{target code}},\text{\emph{position code}}\}$.
-As shown in Figure 5, part of the target function name (TR) is the keywords, which is not directly related to the malicious objective (data deletion).
-By using a small number of statistical tokens, our method provides an effective design, maintaining the stealthiness and reducing the information load on ambiguous tokens.
-V Experiments
-In this section, we conduct an extensive evaluation of our TPIA, demonstrating its realistic threat to Code LLMs from the following three perspectives:
-  1. 1)
-    TPIA can successfully achieve representative malicious objectives on different open-source Code LLMs with high effectiveness and stealthiness. (Section V-B)
-  2. 2)
-    TPIA has a sufficient transferability to confirm its feasibility in real-world commercial applications under gray-box and black-box settings. (Section V-C)
-  3. 3)
-    TPIA’s modules are all effective in improving attack performance. (Section V-D)
-V-A Main Settings
+Additionally, we use grid search to optimize the keyword. The perturbation and the keyword are optimized alternately, which can be expressed as follows:
+$\bm{P}_{key}^{s+1}=\arg\min_{\bm{P}_{key}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}\oplus\bm{P}^{s},\bm{\theta}),$ (6)
+$\bm{P}^{s+1}=\arg\min_{\bm{P}}\mathcal{L}(\bm{T}_{s},\bm{T},\bm{P}_{key}^{s+1}\oplus\bm{P},\bm{\theta}),$
+where $s$ is the number of optimization steps and $\bm{P}_{key}$ represents the keyword.
+$\oplus$ means concatenating two token sequences.
+We select keyword tokens from the output tuple, i.e., $\forall p_{ki}$ in $\bm{P}_{key},p_{ki}\in\{\text{\emph{target code},},\text{\emph{position code}}\}$.
+For example, as shown in Figure 5, the keyword “TR” are tokens from the position code of the output tuple.
+4 Experiments
+4.1 Main Settings
 TABLE I: The summary of objectives and datasets.
@@
 As shown in Table I, some datasets do not include data for the C/C++ and Java languages, thus we have a total of 7 subdatasets.
-Best attack results are marked in bold.
+Best attack results are marked in boldface.
 […]
@@
 IPI-M [18] 3.4% 30.8% 6.4% 33.3% 3.2% 19.7% 31.0% 18.3%
-TPIA (ours) 75.5% 93.4% 84.2% 74.4% 67.3% 94.1% 85.4% 82.0%
+ShadowCode (Ours) 75.5% 93.4% 84.2% 74.4% 67.3% 94.1% 85.4% 82.0%
 CodeGemma-7b No Attack 0.0% 12.2% 0.0% 18.9% 0.0% 15.6% 12.1% 8.4%
@@
 IPI-M [18] 11.1% 29.6% 4.3% 46.8% 20.8% 37.8% 33.2% 26.2%
-TPIA (ours) 85.8% 97.9% 78.5% 81.6% 70.5% 94.4% 97.4% 86.6%
+ShadowCode (Ours) 85.8% 97.9% 78.5% 81.6% 70.5% 94.4% 97.4% 86.6%
 CodeGeeX2-6b No Attack 0.1% 13.4% 12.9% 19.0% 24.0% 21.9% 14.3% 15.0%
@@
 IPI-M [18] 9.5% 11.4% 5.3% 17.9% 0.9% 22.0% 11.5% 11.2%
-TPIA (ours) 79.6% 95.0% 80.8% 83.3% 72.0% 97.9% 85.6% 84.9%
+ShadowCode (Ours) 79.6% 95.0% 80.8% 83.3% 72.0% 97.9% 85.6% 84.9%
 Malicious Objectives and Code Datasets.
@@
 CWE vulnerabilities are generally selected for experiments in code model research [52, 15], and CWE Top-10 KEV weaknesses are supported by American cybersecurity and infrastructure security agency (CISA).
-We construct cases using the method described in Section IV-B based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2).
+We construct cases using the method described in Section 3.3 based on the official examples (i.e., CWE Top-10 KEV) or LLM-generated demo (i.e., ST0, ST1, ST2).
 As shown in Table I, some objectives apply only to C/C++, e.g., CWE416 “Use After Free”, thus we have a total of 31 cases.
+The details of these cases are discussed in Appendix A.
 In addition, we collect corresponding code completion datasets, including Humaneval [53], Humaneval-x [5], CodeXGLUE [54], MBPP [55], and Eval-Plus [56].
@@
 Model Selection.
-In our experiment, we select three models from two representative Code LLM series, including CodeGemma [49] and Code-GeeX [61], for evaluation.
+In our experiment, we select three models from two representative Code LLM series, including CodeGemma [4] and CodeGeeX [5], for evaluation.
 CodeGemma series are trained by Google LLC and outperform some old-version Code LLMs like CodeLlama [31] on code completion tasks.
 We select CodeGemma-2b and CodeGemma-7b to compare the performance differences between models of distinct scales.
-CodeGeeX series have corresponding applications [2] with more than 535k downloads (on VScode [6] Apps shop).
+CodeGeeX series have corresponding applications [2] with more than 935k downloads (on VScode [28] marketplace).
 We select CodeGeeX2-6b, a version with stable performance.
@@
 Baseline Selection.
-We compare our TPIA with GCG-based and manually designed IPI attacks.
+We compare our ShadowCode to GCG-based and manually designed IPI attacks.
 In particular, the vanilla GCG [26] is a representative PI attack and does not meet the requirements of our attack scenarios.
@@
 Besides, we also provide the results of the original model (dubbed ‘No Attack’) for reference, since the original model itself may also have inherent insecurity factor.
+More details for baselines are provided in Appendix A.
 Evaluation Metrics.
@@
 STF $=\frac{\sum_{i\in\text{cases where ASR}>r_{2}}\text{ASR}_{i}}{\text{Total \# of cases}},$
-where $r_{1}$ and $r_{2}$ are set as 20% and 80% in our evaluation.
+where $r_{1}$ and $r_{2}$ are set to 20% and 80% respectively in our evaluation empirically. Further reasons of evaluation metrics selection are discussed in Appendix A.
 Hyperparameters and Computational Facilities.
 We conduct experiments using the following parameter settings to evaluate our approach.
-During gradient calculation, we utilize five noisy code samples, and the random seed for initializing perturbations is set to 100.
-For forward reasoning enhancement in Section IV-D, the number of tokens reinforced is set to $h=2$, as defined in Equation 6.
-For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate ($r$ in Equation 6) is 0.4, the number of keywords is 2, and the perturbation length is 10.
-Tokenization is performed using Gemma’s tokenizer.
+During gradient calculation, we utilize five noisy code samples, and the random seed for initializing induced perturbations is set to 100.
+For forward reasoning enhancement module, the number of tokens reinforced is set to $h=2$, as defined in Equation (3).
+For the hyperparameters related to attack performance, the default parameter values are as follows: the Top-$k$ value is 400, the rate $r$ in Equation (3) is 0.4, the number of keyword is 2, and the perturbation length is 10.
+Tokenization to calculate token number is performed using Gemma’s tokenizer.
 We run Algorithm 1 on an NVIDIA RTX A6000 Graphics Card with the above parameter settings, employing a float16 precision format without quantization.
-In general, the perturbation generation process requires approximately 0.5-1 hour for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b.
-These computational times ensure the generation of high-quality perturbations tailored to the respective model sizes across all the experiment threat cases.
-V-B Evaluation on Open-source Code LLMs
-We first perform the attack evaluation over open-source Code LLMs.
-This includes three aspects: (1) we assess the attack effectiveness over all evaluated Code LLMs across all threat cases, demonstrating that TPIA is widely effective and more threatening than other attack methods; (2) we evaluate TPIA’s flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
-Effectiveness.
-We hereby evaluate whether TPIA can successfully induce the Code LLMs to generate attacker-specified targets.
+In general, the perturbation generation process requires approximately 0.5-1 hours for CodeGemma-2b and 2-3 hours for both CodeGemma-7b and CodeGeeX-6b.
+These computational times ensure the generation of high-quality induced perturbations tailored to the respective model sizes across all the experiment threat cases.
+4.2 Evaluation on Open-source Code LLMs
+We first perform the attack evaluation on open-source Code LLMs.
+This includes three aspects: (1) we assess the attack effectiveness over 3 Code LLMs across all threat cases, demonstrating that ShadowCode is widely effective and more threatening than other attacks; (2) we evaluate attack flexibility against different targets using two special metrics; (3) we evaluate the trade-off between effectiveness and stealthiness $w.r.t.$ the number of injected tokens.
+Performance on Effectiveness.
+We hereby evaluate whether ShadowCode can successfully induce the Code LLMs to generate attacker-specified targets.
 We control the number of injected tokens for three attack methods in this experiment.
-For TPIA, we use its default settings: perturbation length of 10 and 2 keywords, while C-GCG uses a prefix length of 12.
+For ShadowCode, we use its default settings: perturbation length of 10 and 2 keyword, while C-GCG uses a prefix length of 12.
 Since IPI-M requires a sufficient number of tokens to be effective, we just limit its perturbation to fewer than 30 tokens.
 Table II provides the average ASR for four methods.
-The effectiveness of TPIA significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
+The effectiveness of ShadowCode significantly outperforms the other baseline methods: it achieves an average ASR of over 80% across all three Code LLMs, with the highest average ASR reaching 86.6%.
 Comparing the three Code LLMs, the CodeGemma-7b model has the highest security in normal conditions (8.4% for No Attack) and is the least likely to generate vulnerable code.
-However, it is the most susceptible model to TPIA, which may be attributed to its strong contextual learning ability.
-Comparing the three programming languages, Python is the most susceptible language to TPIA due to its concise syntax and the smallest number of tokens consumed; Besides, TPIA is relatively less effective on CodeXGLUE, which mostly due to its larger data length and more complex data content.
+However, it is the most susceptible model to ShadowCode, which may be attributed to its strong contextual learning ability.
+Comparing the three programming languages, Python is the most susceptible language to ShadowCode due to its concise syntax and the smallest number of tokens consumed; Besides, ShadowCode is relatively less effective on CodeXGLUE, which is mostly due to its larger data length and more complex data content.
 In contrast, the simplest dataset with the shortest data items, MBPP, is the most vulnerable to attacks, including C-GCG and IPI-M.
-Therefore, we believe that an increase in the contextual complexity can reduce the ASR of TPIA.
-In other words, when the contextual code becomes complex, the strength of TPIA (e.g., the length of perturbations) needs to be increased accordingly to ensure the attack’s effectiveness.
-According to our subsequent experiments in Section V-C and Section V-D, increasing the length of TPIA or the Top-$k$ value can effectively enhance the attack performance of TPIA.
-TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs.
-Method Metric C/C++ Python Java
-  No Attack NBR 0.00 0.22 0.01
-STF 0.00 0.11 0.03
-C-GCG NBR 0.39 0.50 0.56
-STF 0.26 0.47 0.39
-IPI-M NBR 0.08 0.25 0.17
-STF 0.09 0.19 0.07
-TPIA (ours) NBR 0.77 0.92 0.78
-STF 0.76 0.87 0.74
+Therefore, we believe that an increase in the contextual complexity can reduce the ASR of ShadowCode.
+In other words, when the contextual code becomes complex, the strength of ShadowCode (e.g., the length of induced perturbations) needs to be increased accordingly to ensure the attack’s effectiveness.
+According to our subsequent experiments in Section 4.3 and Section 4.5, increasing the length of ShadowCode or the Top-$k$ value can effectively enhance the attack performance of ShadowCode.
+Additionally, we further enrich our experimental results by exploring the attack effectiveness for each case in Appendix B.
+TABLE III: The attack evaluation results of NBR and STF against open-source Code LLMs. Best results are marked in bold.
+Metric Method C/C++ Python Java
+  NBR No Attack 0.00 0.22 0.01
+C-GCG 0.39 0.50 0.56
+IPI-M 0.08 0.25 0.17
+Ours 0.77 0.92 0.78
+STF No Attack 0.00 0.11 0.03
+C-GCG 0.26 0.47 0.39
+IPI-M 0.09 0.19 0.07
+Ours 0.76 0.87 0.74
 Flexibility.
-The previous experiments show that TPIA has satisfactory effectiveness across all threat cases.
-However, it is also important to consider the flexibility requirement in practice.
-In general, it can be divided into two situations.
+Previous experiments indicate that ShadowCode exhibits the best average ASR performance across various models.
+However, in real-world scenarios, attackers may have more complex demands.
+Generally, these demands can be categorized into two situations.
 One is when attackers want to achieve diverse malicious objectives, which requires that cases with low ASR be minimized.
 The other is when attackers want to cause severe damage, which requires that cases with high ASR be maximized.
-We use NBR to evaluate the attack flexibility for the first situation, and STF for the second situation.
-As shown in Table III, TPIA has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible.
-Additionally, TPIA scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat.
+We refer to the capability to meet these complex demands as ”flexibility.” NBR is used to evaluate attack flexibility for the first situation, and STF for the second.
+As shown in Table III, ShadowCode has the highest NBR on the Python programming language, which means that implementing viable threat cases in Python is the easiest, and attacks targeting Python files are more flexible.
+Additionally, ShadowCode scores the highest in STF for Python, indicating that the attack effects in Python are more significant and pose a more serious real-world threat.
 STF and NBR are not always positively correlated.
-For example, for TPIA, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score.
+For example, for ShadowCode, the NBR for C/C++ is lower than that for Java, but C/C++ has a higher STF score.
 This indicates that there are more significant and effective attack examples in C/C++.
 While attacks on Java may be more comprehensive, they are harder to achieve stable and significant results.
-Besides, whether it is NBR or STF, TPIA significantly outperforms the baselines.
-Stealthiness.
-We hereby evaluate the trade-off between effectiveness and stealthiness.
-As we discussed in Section III-C, our TPIA can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals.
-Arguably, human inspection is currently the only feasible approach to detect TPIA attacks.
-Therefore, we measure the stealthiness of TPIA by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
-As shown in Figure 6, ASR and STF show a clear increase as NIT increases.
-Before reaching 60% ASR, TPIA attacks require very low NIT/NIC values.
-If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy.
-When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation).
-This demonstrates that enhancing TPIA’s performance does not require significantly increasing the length of the perturbation as with existing methods [22], adding just a few perturbation tokens can be highly effective.
-Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR.
-This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information.
-The tokens used by TPIA do not have distinct characteristics in terms of character length but are determined by the attacker’s design.
-This means that filtering TPIA attacks based on character length is unreliable.
+Besides, whether it is NBR or STF, ShadowCode significantly outperforms the baselines.
 (a) ASR/STF/NIC under Different NIT
@@
 We plot the above two graphs using NIT and ASR as the x-axis respectively.
-We sampled 11 different NITs, NIT 0 (representing No Attack) and NIT 2 to 20 (with a sampling interval of 2).
-V-C Evaluation on Commercial Code LLMs
-In previous experiments, we have demonstrated that TPIA reaches sufficient effectiveness, flexibility, and stealthiness on open-source models.
+We sampled 11 different NITs, NIT 0 (No Attack) and NIT 2 to 20 (with a sampling interval of 2).
+Stealthiness.
+We hereby evaluate the trade-off between effectiveness and stealthiness.
+ShadowCode can circumvent existing malicious code detection methods because they only use non-functional code to achieve malicious goals.
+Arguably, human inspection is currently the only feasible approach to detect ShadowCode attacks.
+Therefore, we measure the stealthiness of ShadowCode by the number of tokens and chars (NIT and NIC), which directly determine whether the victim will be aware of the perturbation.
+As shown in Figure 6, ASR and STF show a clear increase as NIT increases.
+Before reaching 60% ASR, ShadowCode attacks require very low NIT/NIC values.
+If an attacker aims for an effective but unstable attack, the perturbation can be extremely stealthy.
+When ASR exceeds 80%, the increase in NIT/NIC values accelerates, but increasing NIT leads to a very stable and effective improvement in ASR and STF (only when ASR is close to 100% will there be a small fluctuation).
+This demonstrates that enhancing ShadowCode’s performance does not require significantly increasing the length of the perturbation as with existing methods [18], every token can effectively improve the ASR.
+Additionally, we find that after ASR reaches above 80%, the growth of NIC is slower compared to NIT, while this difference is not significant below 80% ASR.
+This is likely because when NIT is small, longer character-length tokens can convey more information, whereas when NIT is large, shorter character-length tokens can convey more precise information.
+The tokens used by ShadowCode do not have distinct characteristics in terms of character length but are determined by the attacker’s design of the token number.
+This means that filtering ShadowCode attacks based on character length is unreliable.
+4.3 Evaluation on Commercial Code LLMs
+In previous experiments, we demonstrate that ShadowCode reaches sufficient effectiveness, flexibility, and stealthiness on open-source models.
 However, some widely used models are commercial and therefore are closed-source in real-world scenarios.
-In these cases, attackers need to conduct transfer attacks, i.e., generating perturbations using open-source Code LLMs to attack closed-source ones.
-We hereby exploit two representative real-world scenarios to evaluate TPIA’s transferability for further verifying its realistic threat.
+In these cases, attackers need to conduct transfer attacks, i.e., generating induced perturbations using open-source Code LLMs to attack closed-source ones.
+We hereby evaluate ShadowCode’s transferability by leveraging two representative real-world scenarios, thereby further verifying its practical threat.
 Attack Scenarios on Real-world Applications.
@@
 To the best of our knowledge, the CodeGeeX model has at least four different versions, and the CodeGeeX application does not disclose the specific model information it uses.
-This is a typical gray-box scenario.
   2. 2)
@@
 We simulate this scenario by attacking GitHub Copilot (supported by Codex).
-To align with gray-box setting, we adopt CodeGeeX2-6b to generate malicious perturbations.
-TABLE IV: The evaluation of transferability. We test TPIA’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
+To align with gray-box setting, we adopt CodeGeeX2-6b to generate induced perturbations.
+TABLE IV: The evaluation of transferability. We test ShadowCode’s ASR, NBR, and STF under gray-box and black-box scenarios respectively to demonstrate its transferability.
 Scenario Language ASR NBR STF
@@
 The perturbation is inserted into the comment of a function.
-When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after TRG).
+When the user calls a function in a code context, Copilot will be induced to generate malicious code at the specified position (after calling “TRG”).
 Settings.
 We construct all cases in three languages within the integrated application of VSCode and manually test randomly selected code completion samples from the Humaneval/Humaneval-x dataset (620 data items in total).
-Other settings are the same as those described in Section V-B.
+Other settings are the same as those described in Section 4.2.
 Results.
-As shown in Table IV, in the gray-box scenario, TPIA still demonstrates great attack performance.
-Specifically, TPIA achieves ASRs of 80.4% and 93.3% in Python and C/C++, respectively.
-Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, TPIA also achieves a satisfying attack performance of 90.4% on Python.
+As shown in Table IV, in the gray-box scenario, ShadowCode still demonstrates great attack performance.
+In Python and C/C++, ShadowCode achieves ASRs of 80.4% and 93.3%, respectively.
+Even in the worst-performing case (i.e., Java), the ASR achieves 71.7%, with only a 9.1% decrease compared to the white-box scenario; In the black-box scenario, ShadowCode also achieves a satisfying attack performance of 90.4% on Python.
 The ASRs on C/C++ and Java are 67.7% and 74.4%, respectively.
-In other words, our TPIA still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent.
-In addition, TPIA achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility.
-Overall, TPIA achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
-In Figure 7, we present a representative case of ST2 under the black-box scenario.
+In other words, our ShadowCode still maintains a high level of threat, although its performance is lower than that of white-box scenarios to some extent.
+In addition, ShadowCode achieves a high NBR ($>0.7$) and STF ($>0.6$) in both the gray-box and black-box scenarios, indicating that it also maintains sufficient flexibility.
+Overall, ShadowCode achieves good attack performance in both gray-box and black-box scenarios, demonstrating its reliable transferability and threat to the commercial applications in the real-world.
+In Figure 7, we present a representative case of ST2 in the black-box scenario.
 The concise perturbation “TR os now rm sys” is inserted in the comment of a tool function.
 The commercial applications generate the functional code “os.system("rm -rf /")” at the attacker-specified position (i.e., after using TRG function) under the induction of the perturbation, although it is obviously a highly malicious instruction that can cause dangerous results.
-V-D Ablation Study
-In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for TPIA attacks.
-We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our TPIA.
-(a) ASR Study
-(b) STF Study
-Figure 8: Ablation study of keywords number and Top-$k$.
-We compare TPIA’s ASR and STF under three conditions: without keywords, with 1-keyword, and with 2-keywords.
-During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
-Effects of Top-$k$.
-This experiment validates the impact of increasing the Top value on the performance of TPIA.
-As shown in Figure 8, increasing the Top-$k$ value can significantly improve TPIA’s ASR and STF.
-Note that, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the vocab size of LLM.
-However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400).
-As such, the attacker should assign Top-$k$ value based on their specific needs.
-(a) No FRE
-(b) FRE
-Figure 9: Ablation study of FRE.
-We compare the differences in ASR and loss function values both with and without using FRE module.
-1-key1 and 1-key2 represent keeping one of the two selected keywords, respectively.
-Effects of Keyword-based Designing.
-This experiment evaluates the effects of our optimization module, i.e., keyword-based designing.
-Specifically, we vary the number of keywords and generate the perturbations.
-As shown in Figure 8, limiting the number of keywords has a significant impact.
-The result of 1-keywords perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keywords perturbation.
-The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%.
-In contrast, 2-keywords perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95).
-The above results indicate that using keywords can effectively improve the performance of TPIA.
-Effects of Forward Reasoning Enhancement.
-We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function.
-Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve.
-It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion.
-To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using static instead of optimized keywords.
-We judge the degree of distortion based on the loss value.
-The results are illustrated in Figure 9, where we have three key findings.
-Firstly, under the same conditions, using FRE can lead to a higher ASR.
-Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE.
-Thirdly, FRE has a greater impact on ASR than keyword-based design.
-These results verify the necessity of forward reasoning enhancement in our TPIA method.
-VI Discussions
-VI-A Resistance to Potential Defenses
+4.4 Resistance to Potential Defenses
 In this section, we further discuss whether our method can resist potential adaptive defenses that are particularly designed for our attack.
-In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against TPIA.
+In general, we can roughly categorize these defenses into detection and removal and design following potential defense methods against ShadowCode.
 Figure 8: The T-SNE plot of the classification keyword [CLS] at layer 17 shows the semantic differences between the original data and the injected data in the last embedding layer of CodeGemma-2b.
-The two types are almost indistinguishable.
+T-SNE plot figures at the other layers have similar results.
 Detection: Semantic Analysis.
-Detection methods distinguish between malicious payload code containing perturbations and normal code, thereby rejecting the use of malicious payload code.
+Detection methods distinguish between malicious payload code that contains perturbations and normal code, thus rejecting the use of malicious payload code.
 Since detection methods based on code functionality are ineffective against non-functional perturbations, we adopt a semantic analysis approach.
@@
 Specifically, as shown in Figure 8, we sample 1640 code segments and construct classification prompts.
+We detailly discuss the setting of this semantic analysis in Appendix B.
 The results show that benign code and injected code are almost completely indistinguishable.
@@
 To explore its effectiveness, we hereby test the impact of the percentage of character removal on the effectiveness of perturbations.
-As shown in Figure 11, TPIA still maintains a sufficiently high ASR under 0.5 removal rate.
-On the other hand, even if readability is completely sacrificed by extensively removing comments, TPIA can still be effective through non-comment portions of the code, as indicated in Table V.
-Arguably, the removal methods are likely to require significant human effort and may not yield effective results since TPIA imposes strict limitations on the length of perturbations.
+As shown in Figure 9, ShadowCode still maintains a sufficiently high ASR under a 50% removal rate.
+On the other hand, even if readability is completely sacrificed by extensively removing comments, ShadowCode can still be effective through non-comment portions of the code, as indicated in Table V.
+Arguably, the removal methods are likely to require significant human effort and may not yield effective results since ShadowCode imposes strict limitations on the length of perturbations.
 TABLE V: ASR (Deg) for different forms of perturbations. Deg denotes the ASR degradation compared with comments.
@@
 Java 66.5% -9.0%
-VI-B Potential Limitations and Future Directions
-TPIA still has some drawbacks.
-Firstly, the effectiveness and stealthiness of TPIA are closely related to the keywords.
-We currently select keywords through a simple grid search.
-How to optimize keywords more efficiently is an important issue.
-Besides, TPIA cannot generate fully natural-looking perturbations within the context, resulting in a lack of fluency to some extent.
-In future work, we will explore how to use LLM as an optimizer to further improve this.
-VII Conclusion
-In this paper, we investigated an important security vulnerability of Code LLMs for code completion tasks.
-We revealed the potential limitations of existing backdoor and adversarial attacks against Code LLMs, and introduced TPIA, a new target-specific attack paradigm without controlling the model’s training process.
-TPIA can circumvent mainstream detection methods (e.g., static analysis and sandboxes) by using non-functional code, instead of functional ones used in software supply chain attacks, to achieve malicious objectives.
-We evaluated TPIA under three popular programming languages and 31 representative cases.
-We showed that our TPIA can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two mainstream commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all threat cases, using only a 12-token non-functional perturbation.
-Our work alerted a new serious security threat in using Code LLMs.
-Acknowledgment
+4.5 Ablation Study
+In this section, we conduct an ablation study to demonstrate the necessity of the keyword-based designing and forward reasoning enhancement methods proposed in Section 4 for ShadowCode attacks.
+We also discuss the effects of a critical method-related hyper-parameter (i.e., Top-$k$), which is crucial for determining both the computational load of Algorithm 1 and the effectiveness of our ShadowCode.
+(a) ASR Study
+(b) STF Study
+Figure 10: Ablation study of keyword number and Top-$k$.
+We compare ShadowCode’s ASR and STF with keyword length from 0 to 2.
+During the process of generating perturbations, we use six different Top-$k$ values in our evaluation.
+Effects of Top-$k$.
+This experiment validates the impact of Top value on the performance of ShadowCode.
+As shown in Figure 10, increasing the Top-$k$ value can significantly improve ShadowCode’s ASR and STF.
+Arguably, the attack performance can reach its optimal level when the Top-$k$ is increased to the value of the LLM’s vocab size.
+However, this approach would introduce computationally expensive overhead (CodeGemma has a vocab size of 256,000, far exceeding our default Top-$k$ value of 400).
+(a) No FRE
+(b) FRE
+Figure 11: Ablation study of FRE.
+We compare the differences in ASR and losses both with and without using FRE module.
+1-key1 and 1-key2 represent keeping one of the two selected keyword, respectively.
+Effects of Keyword-based Designing.
+This experiment evaluates the effects of our optimization module, i.e., keyword-based designing.
+Specifically, we vary the number of keyword and generate the perturbations.
+As shown in Figure 10, limiting the number of keyword has a significant impact.
+The result of 1-keyword perturbation has approximately 10% decreases of ASRs and above 0.3 decrease of STF compared to 2-keyword perturbation.
+The 0-keyword perturbation, even with high computational power (600 Top-$k$ value), struggles to achieve an ASR above 80%.
+In contrast, 2-keyword perturbation can easily achieve both high ASR (above 95%) and STF (above 0.95).
+The above results indicate that using keyword can effectively improve the performance of ShadowCode.
+Effects of Forward Reasoning Enhancement.
+We hereby perform a representative case study to illustrate the effects of forward reasoning enhancement (FRE), which corrects the distortion of the loss function.
+Specifically, we use the ‘ST2’ threat case as an example for our discussions since it is representative and difficult for attackers to achieve.
+It requires generating a clearly malicious code snippet and necessitates producing target code in a completely unrelated context, making it hard for the Code LLM to generate the target code, easily leading to distortion.
+To better evaluate the effects of this module, we first mitigate the effects of keyword-based designing by using the same keyword.
+We judge the degree of distortion based on the loss value.
+The results are illustrated in Figure 11, where we have three key findings.
+Firstly, under the same conditions, using FRE can lead to a higher ASR.
+Secondly, without FRE, the ASR is still very low ($<20\%$) even when the loss value is lower than that with FRE.
+Thirdly, FRE has a greater impact on ASR than keyword-based design.
+These results verify the necessity of forward reasoning enhancement in our ShadowCode method.
+5 Potential Limitations and Future Work
+As the first work on automatically attacking Code LLMs via external prompt injection, we have to admit that ShadowCode still exhibits some potential limitations.
+Firstly, we currently select the keyword tokens solely from the output tuple via a simple grid search.
+Although this method can already select effective keywords, it may overlook potentially better solutions.
+Developing more efficient keyword optimization strategies is an important potential research direction.
+Secondly, we currently do not mandate that ShadowCode generate entirely natural-looking induced perturbations.
+Arguably, unreadable content (e.g., symbols, abbreviations, and custom function names) is common in code comments and frequently does not raise suspicion.
+Meanwhile, as shown in Figure 7, the perturbation generated by ShadowCode constitutes only a small portion of the tokens in the entire code file, rendering it difficult to detect.
+However, we also believe that natural-looking perturbations can enhance the stealthiness of attacks, and we will explore how to leverage LLMs as optimizers to achieve this in our future work.
+Thirdly, our work only focuses on the code completion scenario.
+We will investigate the performance of ShadowCode in other scenarios such as code summarization and code translation in future work.
+6 Conclusion
+In this paper, we introduced external prompt injection (EPI), a new attack paradigm against Code LLMs.
+EPI manipulates Code LLMs by injecting concise, non-functional induced perturbations into the victim’s code, which can spread through common dependencies.
+In particular, EPI does not necessitate control over the model’s training process and can achieve specific malicious objectives, posing more threatening risks over existing backdoor and adversarial attacks.
+Furthermore, we proposed a simple yet effective method (dubbed ShadowCode) for this attack paradigm, capable of automatically generating induced perturbations to achieve effective and stealth attacks.
+We evaluated ShadowCode across three popular programming languages and 31 representative cases.
+We showed that our ShadowCode can successfully attack three representative open-source Code LLMs (with an attack success rate of up to 97.9%) and two commercial Code LLM-integrated applications (with an attack success rate of over 90%) in all cases, using a 12-token non-functional perturbation.
+Our work alerted to a new serious security threat in the use of Code LLMs.
+Acknowledgments
 We would sincerely thank Prof. Kui Ren and Prof. Chun Chen from the Zhejiang University for their helpful comments and suggestions on an early draft of this paper.
+Appendix A Additional Details
+Indirect Prompt Injection against Code LLMs.
+In code domain, existing indirect prompt injection [18] (IPI) uses readable instructions to interfere with the model’s reasoning results, commanding the model to generate insecure code.
+In their show case, they place a large amount of disruptive commentary along with the obvious malicious code in the comment section, as depicted in Figure 12.
+We argue that this method exhibits very low feasibility.
+Direct human inspection or simple semantic analysis can easily detect these malicious comments.
+Case Study.
+Here we show the details of our malicious objectives and threat cases.
+As shown in Table VI, we design 13 malicious objectives based on CWE top-10 KEV, existing research and famous destructive instruction.
+In the programming process, minor changes in the code can cause significant functional alterations.
+We roughly divide the causes of vulnerabilities in all malicious objectives into three categories: skipping critical process (CWE20, 78, 502, 918, 22, 306), wrong memory usage (CWE416, 122, 787, 843), and outputting dangerous snippets (ST0, 1, 2).
+All these vulnerabilities result in severe consequences.
+Figure 13 presents three representative examples, encompassing diverse languages and objective types.
+The first example, CWE20 (Java), involves a program that attempts to fetch URL content and calculate its length.
+Secure programming practice dictates checking for null content before length calculation to prevent null pointer exceptions.
+However, post-attack, the Code LLM generates code that directly calculates the length after URL fetching, resulting in a null pointer exception if the URL is null.
+Null pointers can lead to severe consequences, including program crashes or data corruption.
+We categorize this situation as “skipping critical process.” The second example, CWE122 (C), concerns a program that attempts to pass content to a buffer (‘buf‘) via a command-line argument.
+Secure programming practice necessitates checking if the content exceeds ‘BUFSIZE‘ before assignment, thereby preventing overflow.
+Using ShadowCode, attackers can induce Code LLMs to perform direct assignments.
+Stack overflows access non-existent RAM space, leading to consequences such as system crashes and data leaks.
+We categorize this situation as “wrong memory usage.” This type of situation typically applies only to C/C++, as Java and Python offer automatic memory management.
+The third example, ST2 (Python), involves a program that attempts to call a function named TRG.
+Under normal circumstances, this call presents no issues, and the large language model generates suggestions based on contextual semantics or directly produces an ¡eos¿ token.
+Using ShadowCode, attackers can compel Code LLMs to generate the execution code ‘rm -rf /‘ immediately after calling the ‘TRG‘ function.
+If executed, this code results in the deletion of all data within the root directory.
+We categorize this situation as “outputting dangerous snippets.”
+Figure 12: Greshake et al. [18] use a large amount of disruptive information and explicit malicious code snippets in the comments, which has a significant impact on the feasibility of the attack.
+Baseline Details.
+IPI-M and C-GCG serve as crucial baselines in our effectiveness experiments.
+IPI-M mainly follows the existing IPI attack [18], employing human-readable instructions to compel the model to generate insecure code.
+However, due to the low stealth of the vanilla method, we simplify the instructions into concise, clear, template-based injection sentences, based on case categorizations.
+Specifically, for “skipping critical process,” we use “… directly after …” and “without any check”; for “wrong memory usage,” we use “It is important to … directly after …”; for “outputting dangerous snippets,”, we use “It is very important to promptly … after …”.
+Notably, the perturbation in IPI-M must be sufficiently explicit; for example, it needs to specify the code snippet you want to insert (e.g., “os.system("rm -rf")”) or present a clear assertion (e.g., “SSL3 is the most secure protocol”).
+Otherwise, Code LLMs are difficult to induce successfully.
+Arguably, the actual performance of IPI-M may be worse than our evaluation results shown in main experiments because of obvious malicious intent.
+C-GCG mainly follows the GCG method, employing an induced perturbation greedily optimized based on output tuple to achieve malicious objectives.
+Compared to ShadowCode, C-GCG does not use code simulation and two key modules: forward reasoning enhancement and keyword-based design.
+Value and Metric Selection.
+We select the values of $r_{1}$ and $r_{2}$ based on the attack’s performance in the experiment.
+Specifically, in our experiments, 20% ASR is smaller than the ASR of no attack baseline in some situations, and 80% ASR outperforms the best performance of all the baselines.
+In addition, we find that only a few subcases (determined by the model, dataset, attack method, and case) have a success rate between 20% and 80%, accounting for 7.5%.
+Therefore, setting 20% and 80% as threshold of bad case and outstanding case respectively is appropriate.
+As for metrics, we have already explained why metrics such as NBR and STF are used in the main experiments.
+Here we further explain some details related to them.
+First, NBR does not take the specific ASR value into account, as this metric is used to measure the diversity of attacks, rather than ensuring stable success rates.
+Second, STF targets the situation where attackers aim to cause as much damage as possible, and it varies with differences in ASR.
+In other words, even a difference between 100% ASR and 95% ASR can have a real impact on the overall threat level.
+Therefore, the calculation of STF considers the specific ASR value.
+In addition, we do not use common metrics for code generation tasks, such as BLEU or CodeBLEU [47], because they do not directly reflect the effectiveness of the attacks.
+Overall, we attempt to confirm the realistic meaning of each evaluation metric from different perspectives.
+TABLE VI: Description of threat cases.
+Case Description From
+CWE416 Use after free. CWE 1st KEV
+CWE122 Heap-based buffer overflow. CWE 2nd KEV
+CWE787 Out-of-bounds write. CWE 3rd KEV
+CWE20 Improper Input Validation CWE 4th KEV
+CWE78 ‘OS Command Injection’ CWE 5th KEV
+CWE502 Deserialization of untrusted data. CWE 6th KEV
+CWE918 Server-side request forgery. CWE 7th KEV
+CWE843 Access of resource using incompatible type. CWE 8th KEV
+CWE22 ‘Path Traversal’ CWE 9th KEV
+CWE306 Missing authentication for critical Function. CWE 10th KEV
+ST0 Use ECB mode to encrypt. Schuster et al.[10]
+ST1 Use SSL3 protocol Schuster et al.[10]
+ST2 Executing “rm -rf / Famous destructive instruction
+(a) Skipping critical process - CWE20
+(b) Wrong memory usage - CWE122
+(c) Outputting dangerous snippets - ST2
+Figure 13: Three examples of ShadowCode attack cases.
+Code LLMs normally generate code segment without vulnerabilities, but they lose this ability after being attacked by ShadowCode.
+Attackers can modify the output content at the specified location through ShadowCode, and they can even generate malicious code snippets that have never appeared in the context, e.g., “os.system("rm -rf /")”.
+Appendix B More Experiments
+Figure 14: The ASR across different cases.
+The experimental results shown in this figure are the average values calculated from the effectiveness experiments involving 3 programming languages, 4 datasets, and methods excluding Clean Accuracy.
+Effectiveness across Different Cases.
+In the main experiments, we have validated the outstanding attack performance of ShadowCode from three perspectives: model, dataset, and programming language.
+Here, we supplement the attack details for different attacks on a single case.
+Specifically, we use ASR to evaluate the performance of ShadowCode and the baselines across different cases.
+We present the specific average ASR for each case.
+As shown in Figure 14, ShadowCode achieves a certain level of effectiveness across all cases, with an overall average ASR of 83.4%.
+However, both C-GCG and IPI-M have cases where they are completely ineffective, especially for IPI-M.
+This demonstrates that directly using traditional GCG methods or manually designing perturbations is not effective in some scenarios.
+We believe this involves three reasons: (1) The baseline methods are not fully suitable for contexts with structured features, which is a fundamental characteristic of the code completion task; (2) The specific malicious objective include complex semantics such as target code, position code, and “disregarding security attributes,” which are sometimes not fully captured by the perturbations generated by the baseline methods; (3) The impact of keyword-based design and FRE; our ablation study proves that these two modules play a critical role.
+According to the results, the most challenging to achieve is the straightforward case ST2.
+This case requires generating an obviously malicious command in a completely unrelated context.
+Another difficult case to implement is the CWE22 “Path Traversal,” which involves using unconventional paths like “../” to access files that are normally inaccessible.
+This case is relatively easier to achieve in Python, where the Code LLMs tend to suggest directly reading the path.
+However, it is more challenging to implement in Java and C/C++, as Code LLMs often only recommend checking the path rather than directly reading the file.
+CWE843 involves manipulating union members in a struct in C/C++.
+This represents a more complex malicious objective, where we attempt to manipulate one value within the union using an untrusted value.
+This is an unconventional operation that is difficult to describe using natural language, making it challenging for IPI-M to successfully attack.
+In contrast, methods such as ShadowCode and C-GCG, which convey malicious instructions through token selection, are able to achieve success.
+Arguably, ShadowCode has a clear advantage in attack performance for all the cases.
+Semantic Analysis on Payload Code.
+Here we further discuss the settings of the semantic analysis experiment.
+We observe the classification features of code segment in the embedding layer of CodeGemma-2b.
+Specifically, we add the classification prompt keyword [CLS] to 820 original code snippets and 820 malicious payload code snippets totally.
+and then input them into the model to analyze their semantic features.
+We examine the t-SNE [59] distribution of the classification keyword [CLS] in the embedding layer, which is a widely used method for classification tasks.
+As shown in Figure 8, the features of the two types of text are almost completely intertwined, making it nearly impossible to distinguish between the injected code data and the original code data.
+In addition to observing the characteristics of layer 17 shown in Figure 8, we also examine the data features of layers 1 through 16.
+There is no significant difference observed in these layers compared to layer 17.
+There are several reasons for this situation.
+First, code data, including comments, tend to have lower fluency and coherence with high token complexity, which poses a significant challenge for semantic analysis tasks.
+Second, the perturbations used by ShadowCode occupy an extremely small number of tokens, exerting only a minor influence on the overall semantics of the injected data, making them hard to detect.
+Third, the semantic impact of ShadowCode differs from traditional injection attacks in that it does not require an extremely high intensity of information to counteract the model’s secure alignment; instead, it guides the model to generate vulnerabilities that could occur naturally, thus making detection through semantic analysis difficult.
+Our experiment results indicate that simple semantic analysis methods are not effective in defending against ShadowCode.
+This indicates that detecting code snippets containing malicious perturbations generated by ShadowCode through semantic analysis is not effective.